### v2

# Forbidding just one intersection for short integer sequences

Abstract
In this paper, we study the famous Erdős–Sós forbidden intersection problem for words over an alphabet of size $m$: what is the maximal size of a subfamily $\mathcal{F}$ of $m]^{n}$ that does not contain two vectors $x,y$ coinciding on exactly t1t-1 coordinates? We answer this question provided $m\geqslant\operatorname{poly}(t)$ and $n\geqslant\operatorname{poly}(t)$ for some polynomial function $\operatorname{poly}(\cdot)$ of $t$, greatly extending the recent result of Keevash, Lifshitz, Long and Minzer. Our proof combines some of the recently developed methods in extremal combinatorics, including the spread approximation technique of Kupavskii and Zakharov and the hypercontractivity approach developed in a series of works by Keevash, Keller, Lifshitz, Long, Marcus and Minzer.
1 Introduction
Let $\binom{[n]}{k}$ be a family of $k$-subsets of $n]=\{1,\ldots,n\}$. In 1971, Erdős and Sós (see [19]) considered the following problem.
Probem 1.
Fix $n\geqslant k\geqslant t\geqslant 1$. What is the maximal size of a family $\mathcal{F}\subset\binom{[n]}{k}$ such that for any $F_{1},F_{2}\in\mathcal{F}$ we have $|F_{1}\cap F_{2}|\neq t-1$?
Problem 1 is a harder version of the following question raised by Erdős, Ko and Rado [18] ten years earlier.
Probem 2.
Fix $n\geqslant k\geqslant t\geqslant 1$. What is the maximal size of a family $\mathcal{F}\subset\binom{[n]}{k}$ such that for any $F_{1},F_{2}\in\mathcal{F}$ we have $|F_{1}\cap F_{2}|\geqslant t$?
Erdős, Ko and Rado settled it for $n\geqslant n_{0}(k)$ in the famous Erdős–Ko–Rado theorem [18]. While Problem 2 was completely solved by Ahlswede and Khachatrian [1] after several improvements due to Frankl [33], Wilson [63] and Frankl and Füredi [24], Problem 1 remains open. Erdős and Sós conjectured [^1] The original conjecture of Erdős and Sós is for families avoiding intersection size $1$, the general case is due to Erdős. that any extremal solution of Problem 1 has the following form:
$\mathcal{F}_{T}=\left\{F\in\binom{[n]}{k}\;:\;T\subset F\right\}$
for some set $T$ of size $t$, provided $k\geqslant 2t$ and $n\geqslant n_{0}(k)$. [^2] Note that $\mathcal{F}_{T}$ is the unique up to an isomorphism extremal example for Problem 2 if $n\geqslant(t+1)(k-t+1)$ [33, 63]. The conjecture was verified by Frankl and Füredi in the seminal paper [23]. However, the other regimes of parameters, when $t$ and $k$ grow with $n$, are also important and found numerous applications in discrete geometry [28, 39], communication complexity [62] and quantum computing [6]. Great progress on the problem was made by Keller and Lifshitz who solved Problem 1 when $n$ is linear in $k$ [49], i.e. $n\geqslant f(t)k$ for some function $f(\cdot)$. Later, Ellis, Keller and Lifshitz [14] extended the previous result to the regime when $k\leqslant(1/2-\varepsilon)n$ for any small enough $\varepsilon$ provided $n\geqslant n_{0}(t,\varepsilon)$. Recently, Kupavskii and Zakharov [52] managed to solve Problem 1 under the polynomial dependencies between $n,k,t$, e.g. when $n=k^{\alpha}$, $t=k^{\beta}$ with $0<\beta<1/2$ and $\alpha>1+2\beta$, provided $k\geqslant k_{0}(\alpha,\beta)$. Partial results were obtained by Cherkashin [8] and Keevash, Mubayi and Wilson [44]. It is also worth mentioning the celebrated Frankl and Wilson result [31] that bounds the size of $\mathcal{F}$ in Problem 1 by $\binom{n}{k-t}$ provided $k\geqslant 2t-1$ and kt+1k-t+1 is a prime power.
In recent decades, a rich and active direction of research has emerged: extending results from extremal set theory to families of various mathematical objects. The aforementioned Erdős–Ko–Rado theorem [18] has received numerous extensions to families of permutations [21, 7, 56, 12, 52, 46, 55], finite vector spaces [32], integer sequences [2, 29], matchings [40, 34], independent sets of a graph [4, 5, 36, 64, 26](see also the survey [37]) and many other objects [9, 25, 61, 38, 53, 54](see also the survey [17]).
In contrast, there are much fewer works on generalization of Problem 1 to other domains, mostly due to its hardness even in the $\binom{[n]}{k}$ case. There are some results for permutations [16, 15, 43, 48], vector spaces [11] and integer sequences [42]. Significant progress in this direction was made by Kupavskii and Zakharov [52]. In particular, they present results for subfamilies of $m]^{n},\binom{[n]}{k}^{w}$, permutations and $(n,k,w)$-designs. However, assumptions of their theory turn out to be quite restrictive in our setup, which we discuss next.
In this paper, we focus on Problem 1 adapted to subfamilies of $m]^{n}$ for some integers $m,n$. For two vectors $x,y\in[m]^{n}$, we define their agreement $\operatorname{agr}(x,y)$ by
$\operatorname{agr}(x,y)=|\{i\in[n]\;:\;x_{i}=y_{i}\}|.$
Then, the question we study in this paper is as follows.
Probem 3.
Let $m,n,t$ be integers such that $m\geqslant 2$ and $n\geqslant t\geqslant 1$. What is the maximal size of a subfamily $\mathcal{F}\subset[m]^{n}$ such that for any $x,y\in\mathcal{F}$ we have $\operatorname{agr}(x,y)\neq t-1$?
We will refer to such families $\mathcal{F}$ as (t1)(t-1)-avoiding.
Note that for $t=1$ this problem was resolved independently by Frankl and Tokushige [30] and Ahlswede and Khachatrian [2]. A more general result was obtained by Keevash and coauthors [42].
Theorem 1 (Theorem 1.1 of [42]).
Define families
$\mathcal{S}_{t,r}[m]^{n}=\{x\in[m]^{n}\;:\;|\{j\in[t+2r]\;:\;x_{j}=1\}|\geqslant t+r\}.$
Then, for any $t\in\mathbb{N}$ there is $n_{0}$ such that if $\mathcal{F}\subset[m]^{n}$ is a (t1)(t-1)-avoiding family with $m\geqslant 3$ and $n\geqslant n_{0}$, then $|\mathcal{F}|\leqslant\max_{r\geqslant 0}|\mathcal{S}_{t,r}[m]^{n}|$ with equality only if $\mathcal{F}$ is isomorphic to $\mathcal{S}_{t,r}[m]^{n}$ for some $r\geqslant 0$.
The function $n_{0}(t)$ depends on $t$ super-exponentially in their proof.
Observe that when $m>t$ the unique maximum in $\max_{r}|\mathcal{S}_{t,r}[m]^{n}|$ is attained for $r=0$. Thus, if $m>t$, then the only extremal families are $t$-stars $\{x\in[m]^{n}\;:\;x_{T}=z\}$ for some $T\subset[n]$ of size $t$ and $z\in[m]^{T}$, provided $n\geqslant n_{0}(t)$.
Another result was obtained by Kupavskii and Zakharov [52].
Theorem 2 (Proposition 23 of [52]).
Suppose that integers $m,n,t$ satisfy
$m\geqslant\max\{2^{15}\ln n,2^{17}q\ln q\}\text{ and }n\geqslant 2q$
for $q=800t^{2}\ln m$. If $\mathcal{F}\subset[m]^{n}$ is a (t1)(t-1)-avoiding family and $|\mathcal{F}|\geqslant 2/3\cdot m^{n-t}$, then there exists a $t$-star $\mathcal{S}_{(T,z)}=\{x\in[m]^{n}\;:\;x_{T}=z\}$ for some $T\subset[n]$ of size $t$ and $z\in[m]^{T}$ such that $|\mathcal{F}\setminus\mathcal{S}_{(T,z)}|\leqslant m^{-4t}|\mathcal{S}_{(T,z)}|$.
Despite the restrictions $m\geqslant 2^{15}\ln n$ and $n\geqslant 1600t^{2}\ln m$, their finding suggests that the result of Theorem 1 can be potentially obtained when $m,n$ depend polynomially on $t$. Thus, it is natural to conjecture that the conclusion of Theorem 1 holds when $m\geqslant\operatorname{poly}(t)$ and $n\geqslant\operatorname{poly}(t)$, without logarithmic lower bounds on $m$ and $n$. Here $\operatorname{poly}(\cdot)$ hides some polynomial functions. Our main result confirms this conjecture.
Theorem 3.
There exist polynomials $n_{0}(\cdot)$ and $m_{0}(\cdot)$ such that for all $t\geqslant 1,n\geqslant n_{0}(t),m\geqslant m_{0}(t)$ the following holds. If $\mathcal{F}\subset[m]^{n}$ is (t1)(t-1)-avoiding, then $|\mathcal{F}|\leqslant m^{n-t}$ and the equality holds if and only if $\mathcal{F}=\mathcal{S}_{(T,z)}$ for some $T\subset[n]$ of size $t$ and $z\in[m]^{T}$.
Remark. One can choose $m_{0}(t)=O(t^{300})$. The polynomial $n_{0}(t)$ can be replaced by a function $\tilde{n}_{0}(t)=Ct^{3}\ln t$ for some absolute constant $C$.
The proof combines ideas from Keevash and coauthors [42], Kupavskii and Zakharov [52], some recent developments of Kupavskii [54], a paper of Keller and coauthors on hypercontractivity for some pseudorandom functions [45] and a paper of Kupavskii and the second author [50].
Our proof technique can be adapted to Problem 1, yielding the solution when $k\geqslant\operatorname{poly}(t)$ and $n\geqslant\operatorname{poly}(t)\cdot k$, but we postpone it to more general theory [51].
The paper is organized as follows. In Section 2, we introduce the necessary notation. In Section 3, we present the main tools used in the proof. In Section 4, we develop some theory necessary for the analysis of the case $m\geqslant\operatorname{poly}(t)$ and $n\geqslant\operatorname{poly}(t)\ln m$. In Section 5, we present the solution of Problem 3 when $m\geqslant\operatorname{poly}(t)$ and $n\geqslant\operatorname{poly}(t)\ln m$. In Section 6, we present the solution of Problem 3 when $m\geqslant n^{\operatorname{poly}(t)}$ and $n\geqslant\operatorname{poly}(t)$. Finally, in Section 7, we combine the results of Sections 5 and 6 to prove Theorem 3.
2 Notation
We use standard notation $[n]$ for the set $\{1,\ldots,n\}$ and $X\choose k$} for the set of all $k$-element subsets of $X$. Given families $\mathcal{F}$, $\mathcal{S}$ and a set $X$, we define
$\mathcal{F}(X):=\{A\setminus X:\ A\in\mathcal{F},\,X\subset A\}$
$\mathcal{F}[X]:=\{A:\ A\in\mathcal{F},\,X\subset A\}$
$\mathcal{F}[\mathcal{S}]:=\bigcup\limits_{X\in\mathcal{S}}\mathcal{F}[X].$
$\mathcal{F}(\mathcal{S}):=\bigcup\limits_{X\in\mathcal{S}}\mathcal{F}(X).$
We denote $\partial_{t}\mathcal{F}$ the set of all $t$-shadows of $\mathcal{F}$, i.e. all $t$-element subsets of sets from $\mathcal{F}$. We also denote the collection of all shadows of size at most $t$ as
$\partial_{\leqslant t}\mathcal{F}=\bigcup_{0\leqslant q\leqslant t}\partial_{q}\mathcal{F}.$
Lastly, by $\mathcal{F}\vee\{T\}$ we mean $\{F\cup T\;:\;F\in\mathcal{F}\}$.
We say that two sets $F_{1},F_{2}$ form a matching, if $F_{1}\cap F_{2}=\varnothing$. Two families of sets $\mathcal{F}_{1},\mathcal{F}_{2}$ cross-contain a matching (or contain a cross-matching) if there are two sets $F_{1}\in\mathcal{F}_{1}$, $F_{2}\in\mathcal{F}_{2}$ such that $F_{1}\cap F_{2}=\varnothing$.
Depending on the context, we will identify the families of codes $\mathcal{F}\subset[m]^{n}$ with families of sets from some ambient family $\mathcal{C}\subset\binom{I_{1}\sqcup I_{2}\ldots\sqcup I_{n}}{n}$, where $I_{i}$ consists of integers from (i1)m+1(i-1)m+1 to $im$. We identify each code $x\in[m]^{n}$ with a set $X$ that intersects each $I_{i}$ by the element $(i-1)m+x_{i}$ only. Then $\mathcal{C}$ is a family of all such sets $X$:
$\mathcal{C}=\left\{X\in\binom{[mn]}{n}\;:\;\text{for all }i\in[n]\text{ we have }|X\cap I_{i}|=1\right\}.$ (1)
Given $Z\subset[n]$ and $x\in[m]^{T}$, we call $x$, or, if we want to emphasize the set of coordinates, $(Z,x)$, a partial code or a restriction. Partial codes are shadows of elements of $m]^{n}$. We identify them with shadows of elements of $\mathcal{C}$ in the following way: a restriction $(Z,x)$ corresponds to
$T=\{i(z-1)+x_{z}\;:\;z\in Z\}\in\partial_{|Z|}\mathcal{C}.$
We write
$T=(Z,x);\quad\mathcal{F}[Z\to x]=\mathcal{F}[T];\quad\mathcal{F}(Z\to x)=\mathcal{F}(T).$
If $|Z|=1,$ i.e. $Z=\{i\}$ for some $i\in[n]$, we may write $\mathcal{F}[i\to x]$ and $\mathcal{F}(i\to x)$ instead of $\mathcal{F}[\{i\}\to x]$ and $\mathcal{F}(\{i\}\to x)$.
Given an integer $t$, a set $T$ of size $t$, and a code $z\in[m]^{T}$, we refer to a family $\{x\in[m]^{n}\;:\;x_{T}=z\}$ as $t$-star. We call subfamilies of $t$-stars incomplete $t$-stars.
Sometimes we consider a restriction of a code or a family of codes on a particular set of coordinates. Given $Z\subset Z_{1}\subset[n]$ and $y\in[m]^{Z_{1}}$, by $y|_{Z}$ we mean $y^{\prime}\in[m]^{Z}$ such that $y_{z}=y^{\prime}_{z}$ for all $z\in Z$. For a family $\mathcal{F}$ we denote its restriction $\mathcal{F}|_{Z}=\{y|_{Z}\;:\;y\in\mathcal{F}\}$. We are also interested in the agreement of two partial codes on a subset of the set of their common coordinates. Specifically, given sets $_{1},Z_{2}\subset[n], ZZ$$\subset Z_{1}\cap Z_{2}$ and partial codes $x_{1}\in[m]^{Z_{1}}$, $x_{1}\in[m]^{Z_{2}}$, we denote
$\operatorname{agr}_{Z}(x_{1},x_{2})=\operatorname{agr}(x_{1}|_{Z},x_{2}|_{Z}).$
Finally, we discuss the notation concerning measures and their restrictions. We consider probability measures on $m]^{n}$ or $m]^{Z},Z\subset[n]$. Two common examples are the uniform measure $\mu$ and the class of product measures, i.e., the measures $\nu$ that could be represented as $\nu(x)=\prod_{i\in Z}\nu_{i}(x_{i})$, where $\nu_{i}$ are some probability measures on $[n]$. Note that a uniform measure is also a product measure although we mention it separately. Given a partial code $(Z,x)$ and a measure $\nu$ on $m]^{n}$, the restricted measure $\nu_{Z\to x}$ on $m]^{n}(Z\to x)$ is defined as
$\nu_{Z\to x}(y)=\frac{\nu((x,y))}{\nu([m]^{n}[Z\to x])},$
where $(x,y)$ is a code that coincides with $x$ on $Z$ and with $y$ on $n]\setminus Z$. If $Z=\{i\}$, we sometimes write $\nu_{i\to x}$ instead of $\nu_{\{i\}\to x}$.
For a product measure $\nu=\bigotimes_{i=1}^{n}\nu_{i}$, we define its restriction on the set of coordinates $Z$ as
$\nu_{Z}=\prod_{i\in Z}\nu_{i}.$
Note that, for a product measure, $\nu_{Z\to x}$ does not depend on $x$, and we have $\nu_{Z\to x}=\nu_{[n]\setminus Z}$.
3 Preliminaries
We treat the regimes $\operatorname{poly}(t)\leqslant m\leqslant e^{Cn/t^{2}}$ and $m\geqslant n^{Ct}$ separately, and each regime requires a completely different set of tools. Note that for our assumptions on $n$ these regimes overlap.
We treat the case $m\leqslant e^{Cn/t^{2}}$ using a measure boosting argument, which dates back to Dinur and Friedgut [10]. We illustrate it on subfamilies of sets from some ambient family $\mathcal{A}$ (which is $\mathcal{C}$ in our case) endowed with the uniform measure $\mu(\cdot)$. The argument is based on the following informal observation: given two pseudorandom families $\mathcal{F}_{1},\mathcal{F}_{2}$, under mild assumptions on their measure $\mu(\cdot)$ in some ambient family $\mathcal{A}$, one can obtain families $\widetilde{\mathcal{F}}_{1},\widetilde{\mathcal{F}}_{2}$ in (probably, different) ambient family $\widetilde{\mathcal{A}}$ such that
  1. 1.
[…]
  2. 2.
    if $\widetilde{\mathcal{F}}_{1},\widetilde{\mathcal{F}}_{2}$ cross-contain a matching, then $\mathcal{F}_{1},\mathcal{F}_{2}$ do.
The final trick of this argument is that one can easily prove the existence of cross-matching in families of large measure, leading to the conclusion: pseudorandom families cross-contain a matching. It turns out that, in many Turán-type problems, extremal families admit decomposition into relatively simple pseudorandom subfamilies, and the above reasoning allows to relax conditions on problem parameters.
The efficiency of this idea significantly varies with the notion of pseudorandomness and the range family $\widetilde{\mathcal{A}}$ used. The original definition of pseudorandomness due to Dinur and Friedgut called uncapturability plays the key role in the celebrated junta method [49, 11, 42, 57]. Later, Alweiss and coathours [3] made a breakthrough result in the Erdős sunflower problem using another notion of pseudorandomness called spreadness. Using their results together with delicate combinatorial arguments, Kupavskii and Zakharov [52] developed the spread approximation technique for which they introduced a new concept of pseudorandomness named $\tau$-homogeneity. They apply their method to obtain solutions of the Deza–Frankl problem on $t$-intersecting permutations and the Erdős–Sós forbidden intersection problem under polynomial dependencies between parameters. The technique was further improved in subsequent works of Kupavskii and coathours [53, 54, 55, 50, 27, 38] yielding a large variety of results in extremal set theory. In particular, in this paper we enhance the spread approximation technique for the families of codes. At the same time, Keevash and coathours [41, 42] presented the notion of global families, and using it they reached large progress in Turán-type problems on forbidden large sparse hypergraphs. Then, independently of Kupavskii and Zakharov, they modified their definition of global families (so it coincides now with the $\tau$-homogeneity), and generalized it for functions on a product space [45]. The main result of their findings is a new hypercontractivity inequality for global functions, which we use below; see Theorem 10. In follow-up papers, they show that their approach has a lot of consequences in different fields of discrete mathematics [35, 58, 47, 20].
[…]
3.1 Random gluings
We start with some definitions.
Definition 1 (Definition 6.5 [42]).
We refer to a map $\pi:[m_{1}]\to[m_{2}]$ as a gluing in the sense, that it merges elements of $\pi^{-1}(y),y\in[m_{2}]$, into one element $y$. A gluing $\pi$ is called $b$-balanced, if
$|\pi^{-1}(y)|\leqslant\frac{bm_{1}}{m_{2}}.$
The family of $b$-balanced gluings from $m_{1}]$ to $m_{2}]$ is denoted by $\Pi_{m_{1},m_{2},b}$. If $m_{2}\,|\,m_{1}$ and $b=1$, we omit subscript $b$.
A $b$-balanced gluing $\pi:[m_{1}]^{n}\to[m_{2}]^{n}$ is defined as a tensor product of $b$-balanced gluings $\pi_{1},\ldots,\pi_{n}$, i.e. $\pi(x_{1},\ldots,x_{n})=(\pi_{1}(x_{1}),\ldots,\pi_{n}(x_{n}))$. The set of $b$-balanced gluings from $m_{1}]^{n}$ to $m_{2}]^{n}$ by $\Pi_{m_{1},m_{2},b}^{\otimes n}$. As before, the subscript $b$ is omitted if $b=1$ and $m_{2}\,|\,m_{1}$.
Then, for a family $\mathcal{F}\subset[m_{1}]^{n}$ and a gluing $\pi$, we define
$\mathcal{F}^{\pi}=\{\pi(x)\;:\;x\in\mathcal{F}\}.$
If $\mathcal{F}\subset[m]^{S}$ for some $S\subset[n]$, then, slightly abusing notation, we denote $\mathcal{F}^{\pi|_{S}}$ by $\mathcal{F}^{\pi}$ as well.
Note that if $\mathcal{F}_{1}^{\pi},\mathcal{F}_{2}^{\pi}$ cross-contain a disagreement, then $\mathcal{F}_{1},\mathcal{F}_{2}$ do. Similarly, for a product measure $\nu=\bigotimes_{i=1}^{n}\nu_{i}$ on $m_{1}]^{n}$ and a gluing $\pi=\bigotimes_{i=1}^{n}\pi_{i}\in\Pi_{m_{1},m_{2},b}^{\otimes n}$, we introduce a measure $\nu^{\pi}(\cdot)$ on $m_{2}]^{n}$, defined as follows:
$\nu^{\pi}(\{y\})=\prod_{i=1}^{n}\nu(\pi^{-1}_{i}(y_{i})).$
Let us comment now on the $b$-balanced property of a random gluing. In what follows, it will be important to ensure that if a measure $\nu$ is similar to a uniform measure in some sense, then the measure $\nu^{\pi}(\cdot)$ also is. We control this similarity introducing the notion of $b$-balanced measures.
Definition 2 (Definition 6.7 [42]).
A product measure $\nu=\bigotimes_{i=1}^{n}\nu_{i}$ on $m]^{n}$ is called $b$-balanced if
$\nu_{i}(\{x\})\leqslant\frac{b}{m}$
for any $x\in[m]$ and $i\in[n]$.
One can easily verify the following claim.
Claim 4.
If a product measure $\nu(\cdot)$ on $m_{1}]^{n}$ is $b_{1}$-balanced and a gluing $\pi:[m_{1}]^{n}\to[m_{2}]^{n}$ is $b_{2}$-balanced, then $\nu^{\pi}$ is $b_{1}b_{2}$-balanced.
3.1.1 Gluings and the measure boosting argument
[…]
Claim 5 (Claim 6.8 [42]).
For any $\mathcal{F}\subset[m_{1}]^{n}$, a gluing $\pi:[m_{1}]^{n}\to[m_{2}]^{n}$ and product measure $\nu$ on $m_{1}]^{n}$, we have $\nu^{\pi}(\mathcal{F}^{\pi})\geqslant\nu(\mathcal{F})$.
Clearly, Claim 5 is insufficient for upgrading the measure of a family $\mathcal{F}\subset[m]^{n}$. The following lemma shows how one can enlarge the measure of a family $\mathcal{F}$ choosing a gluing $\pi$ via some averaging argument. The statement of the lemma requires some basic knowledge concerning Markov chains; here we provide some definitions.
For a measure $\nu$ on $m]^{n}$, define the scalar product $\langle\cdot,\cdot\rangle_{\nu}$ on the space $L_{2}([m]^{n},\nu)$ as
$\langle f,g\rangle_{\nu}=\mathbb{E}_{\mathbf{x}\sim\nu}f(\mathbf{x})g(\mathbf{x}).$
Then, let $T$ be a Markov chain with the stationary measure $\nu$. We define a linear operator $T:L_{2}([m]^{n},\nu)\to L_{2}([m]^{n},\nu)$ as follows. Let $\mathbf{y}$ be a step of the Markov chain $T$ starting from a point $x$. Then, we have
$(Tf)(x)=\mathbb{E}f(\mathbf{y}).$
Finally, for a Markov chain $T$ with the stationary measure $\nu$ define $\lambda_{*}(T)$ as follows
$(1-\lambda_{*}(T))^{2}=\sup\{\mathbb{E}_{\mathbf{x}\sim\nu}(Tf(\mathbf{x}))^{2}\;:\;\mathbb{E}_{\mathbf{x}\sim\nu}f(\mathbf{x})=0,\;\mathbb{E}_{\mathbf{x}\sim f}f^{2}(\mathbf{x})=1\}.$
Then, we are ready to state the lemma.
Lemma 6.
Let $\nu(\cdot)$ be a product measure on $m_{1}]^{n}$. Fix $m_{2},b$ such that $m_{2}\leqslant m_{1}\leqslant bm_{2}$. Then, there exists a product Markov chain $T=\otimes_{i=1}^{n}T_{i}$ with the stationary measure $\nu(\cdot)$, such that
$\mathbb{E}_{\bm{\pi}\sim U(\Pi_{m_{1},m_{2},b}^{\otimes n})}\nu^{\bm{\pi}}(\mathcal{F})\geqslant\frac{\nu^{2}(\mathcal{F})}{\langle\mathds{1}_{\mathcal{F}},T(\mathds{1}_{\mathcal{F}})\rangle_{\nu}},$ (2)
where $U(\Pi_{m_{1},m_{2},b}^{\otimes n})$ stands for the uniform measure on $\Pi_{m_{1},m_{2},b}^{\otimes n}$, $\mathds{1}_{\mathcal{F}}$ stands for the characteristic function of $\mathcal{F}$. Moreover, we have $\lambda_{*}(T)\geqslant 1/6$.
For the proof, we refer the reader to that of Lemma 6.9 [42].
It will be convenient to upper bound the scalar product in the right-hand side of (2), replacing the Markov chain operator $T$ with some operator defined as follows. Given a vector $x\in[m]^{n}$ and a product measure $\nu=\bigotimes_{i=1}^{n}\nu_{i}$, define a distribution $N_{\rho}(x)=\otimes_{i=1}^{n}N_{\rho}(x_{i})$, where $N_{\rho}(x_{i})$ is a distribution of the following random variable $\mathbf{y}_{i}$:
[…]
Then, we define an operator $T_{\rho}:L_{2}([m]^{n},\nu)\to L_{2}([m]^{n},\nu)$ by the averaging a function $f$ over $\mathbf{y}$:
$(T_{\rho}f)(x)=\mathbb{E}_{\mathbf{y}\sim N_{\rho}(x)}f(\mathbf{y}).$
We also introduce a quantity that is closely related to the noise operator. Namely, we define the stability of a function $f$ or a family $\mathcal{F}$ as
$\operatorname{Stab}_{\rho}(f)=\langle f,T_{\rho}f\rangle_{\nu}\quad\text{and}\quad\operatorname{Stab}_{\rho}(\mathcal{F})=\langle\mathds{1}_{\mathcal{F}},T_{\rho}\mathds{1}_{\mathcal{F}}\rangle_{\nu}.$
Then, we upper bound $\langle f,Tf\rangle_{\nu}$ by the following lemma.
Lemma 7 (Lemma 5.13[42]).
Suppose that a product Markov chain $T$ has $\lambda_{*}(T)\geqslant 1/6$. Then, we have
$\langle f,Tf\rangle_{\nu}\leqslant\operatorname{Stab}_{5/6}(f).$
Lemma 7 and Lemma 6 yield the following corollary.
Corollary 8.
Fix numbers $n,m_{1},m_{2},b$ such that $m_{2}\leqslant m_{1}\leqslant bm_{2}$. Then, for a function $f\in L_{2}([m_{1}]^{n},\nu)$, a family $\mathcal{F}$ and a product measure $\nu$, there exists a gluing $\pi\in\Pi_{m_{1},m_{2},b}^{\otimes n}$ such that
$\nu^{\pi}(\mathcal{F}^{\pi})\geqslant\frac{\nu^{2}(\mathcal{F})}{\operatorname{Stab}_{5/6}(\mathcal{F})}.$
Due to Corollary 8, for the measure boosting argument, we need to control $\operatorname{Stab}_{5/6}(\mathcal{F})$. The following proposition ensures that it is enough to upper bound $\operatorname{Stab}_{\rho_{0}}(\mathcal{F})$ for sufficiently small $\rho_{0}$.
Proposition 9 (Lemma 5.11 [42]).
For any function $f\in L_{2}([m]^{n},\nu)$, we have
$\operatorname{Stab}_{\rho}(f)\leqslant\|f\|_{2}^{2(1-1/t)}\operatorname{Stab}_{\rho^{t}}(f)^{1/t}$
whenever $t=2^{d}$ with $d\in\mathbb{N}$.
To control $\operatorname{Stab}_{\rho_{0}}(\mathcal{F})$, one need to guarantee some pseudorandomness of $\mathcal{F}$. Authors of the original paper [42] used the notion of globalness due to Keevash et al. [41]. Instead, we use the concept of globalness from [45].
Theorem 10 (Theorem 1.4 from [45]).
Say that a function $f$ on a product space $\prod_{i=1}^{n}\Omega_{i}$ endowed with a product measure $\nu=\otimes_{i=1}^{n}\nu_{i}$ is $r$-global if for any $S\subset[n]$ and any $x\in\prod_{i\in S}\Omega_{i}$, we have
$\|f_{S\to x}\|_{2}\leqslant r^{|S|}\|f\|_{2},$
where $f_{S\to x}:\prod_{i\in[n]\setminus S}\Omega_{i}\to\mathbb{R}$ stands for the function $f_{S\to x}(y)=f(x,y)$. Let $q\geqslant 2$. If $r\geqslant 1$ and $\rho\leqslant\frac{\ln q}{32rq}$, then $\|T_{\rho}f\|_{q}\leqslant\|f\|_{2}$ for any $r$-global function $f$.
In the paper [45], the proof of Theorem 10 is presented for the case when all $\Omega_{i}$ and $\nu_{i}$ coincide. However, it works for the product space since it based on the freezing argument and induction.
We say that a family $\mathcal{F}$ is $\tau$-homogeneous if $\mathds{1}_{\mathcal{F}}$ is $\sqrt{\tau}$-global. This yields the following definition.
Definition 3.
Say that $\mathcal{F}\subset[m]^{n}$ is $\tau$-homogeneous w.r.t. to a product measure $\nu$ if for any $S$ and $x\in[m]^{S}$ we have
$\nu_{S\to x}(\mathcal{F}(S\to x))\leqslant\tau^{|S|}\nu(\mathcal{F}).$
We discuss $\tau$-homogeneous families in more details in the subsequent section. For now, let us present how to bound $\operatorname{Stab}_{\rho_{0}}(\mathcal{F})$ for $\tau$-homogeneous family $\mathcal{F}$. By the Hölder inequality, we have
$\operatorname{Stab}_{\rho_{0}}(\mathcal{F})\leqslant\|T_{\rho_{0}}\mathds{1}_{\mathcal{F}}\|_{4}\|\mathds{1}_{\mathcal{F}}\|_{4/3}.$
If $\rho_{0}$ satisfies the conditions of Theorem 10 for $q=4$, we have
$\operatorname{Stab}_{\rho_{0}}(\mathcal{F})\leqslant\|\mathds{1}_{\mathcal{F}}\|_{2}\|\mathds{1}_{\mathcal{F}}\|_{4/3}=\nu^{5/4}(\mathcal{F}).$
It turns out, that this bound is sufficient for our purposes, see the proof of Lemma 25.
Given two homogeneous families $\mathcal{F}_{1},\mathcal{F}_{2}$, suppose for a while that, iteratively using Corollary 8, we found families $\widetilde{\mathcal{F}}_{1},\widetilde{\mathcal{F}}_{2}\subset[m^{\prime}]^{n^{\prime}}$ such that $\widetilde{\mu}(\widetilde{\mathcal{F}}_{1}),\widetilde{\mu}(\widetilde{\mathcal{F}}_{2})$ is larger than some constant and $\widetilde{\mu}$ is an enough balanced product measure. Then, the following lemma allows us to find a cross-disagreement in families $\widetilde{\mathcal{F}}_{1},\widetilde{\mathcal{F}}_{2}$.
Lemma 11 (Lemma 5.9 [42]).
Let $\nu=\bigotimes_{i=1}^{n}\nu_{i}$ be a product probability measure on $m]^{n}$ such that $\nu_{i}(x)\leqslant\lambda\leqslant 1/2$ for all $i\in[n]$ and $x\in[m]$. Suppose that $\mathcal{G}_{1},\mathcal{G}_{2}\subset[m]^{n}$ are cross-intersecting with $\nu(\mathcal{G}_{i})=\alpha_{i}$ for $i=1,2$. Then
$\alpha_{1}\alpha_{2}\leqslant\left(\frac{\lambda}{1-\lambda}\right)^{2}(1-\alpha_{1})(1-\alpha_{2}).$
There is an alternative measure boosting argument that performs well if one of the initial measures $\mu(\mathcal{F}_{1}),\mu(\mathcal{F}_{2})$ is large. We present it as the following lemma. While its statement can not be found in paper [42], the proof coincides with that of Theorem 2.3 [42]. For the sake of completeness, we provide the proof in the appendix.
Lemma 12.
Let $\mathcal{F}_{1},\mathcal{F}_{2}\subset[m]^{n}$ and let $\mu$ be uniform measure on $m]^{n}$. If $\mu(\mathcal{F}_{1})+\mu(\mathcal{F}_{2})^{\log_{m}(2)}>1$, then there exist $x_{i}\in\mathcal{F}_{i}$ such that $\operatorname{agr}(x_{1},x_{2})=0$.
In the next section, we move on to the following questions: how to decompose a family $\mathcal{F}$ into homogeneous families and how to reduce a question about avoiding agreement of size exactly t1t-1 into a problem of finding a cross-disagreement in homogeneous families?
3.2 Homogeneous families
First, we discuss how to decompose a family $\mathcal{F}$ into homogeneous families $\mathcal{F}_{i}$ up to some small remainder. We start with the following observation.
Lemma 13.
Let $\mathcal{F}\subset[m]^{n}$. Let $\nu$ be a product measure on $m]^{n}$ and fix $\tau>1$. If $Z\subset[n]$ is maximal such that for some $r\in[m]^{Z}$ we have $\nu_{Z\to r}(\mathcal{F}(Z\to r))\geqslant\tau^{|Z|}\nu(\mathcal{F})$, then $\mathcal{F}(Z\to r)$ is $\tau$-homogenous with respect to $\nu_{Z\to r}$. Note that if $Z=[n]$, then $\nu_{Z\to r}(\mathcal{F}(Z\to r))=\mathds{1}\{r\in\mathcal{F}\}$.
Using the above lemma iteratively, one can obtain the decomposition which is called the spread approximation of $\mathcal{F}$.
Lemma 14 (Lemma 52 from [50]).
Fix an integer $q$ and a real number $\tau$. Let $\mathcal{F}\subset[m]^{n}$ and let $\mu$ be the uniform measure on $m]^{n}$. There exists a system of pairs $\{(Z_{i},x_{i})|i=1,\ldots,l\}$ such that for all $i$ we have $Z_{i}\subset[n],|Z_{i}|\leqslant q$, $x_{i}\in[m]^{Z_{i}}$ and a partition of $\mathcal{F}$ into families $\mathcal{F}_{i}\subset\mathcal{F}[Z_{i}\rightarrow x_{i}]$ and $\mathcal{R}$ such that
$\mathcal{F}=\mathcal{R}\sqcup\bigsqcup_{i=1}^{l}\mathcal{F}_{i},$
$\mu(\mathcal{R})\leqslant\tau^{-q}$ and for all $i$ we have $\mu_{Z_{i}\rightarrow x_{i}}(\mathcal{F}_{i}(Z_{i}\rightarrow x_{i}))\geqslant\tau^{-q}$ and $\mathcal{F}_{i}(Z_{i}\rightarrow x_{i})$ is $\tau$-homogenous with respect to $\mu_{Z_{i}\rightarrow x_{i}}$.
In the paper [50], the authors require $\mu(\mathcal{R})\leqslant 32\tau^{-q}$, but their proof works with the tighter bound suggested in the above lemma. For the sake of completeness, let us present the proof of Lemma 14.
Proof of Lemma 14.
We will obtain the decomposition into families $\mathcal{R},\mathcal{F}_{1},\ldots,\mathcal{F}_{l}$ at the end of the following procedure. Define $\mathcal{G}_{0}=\mathcal{F}$. In the step $i$, $i=1,2,\ldots$, find a maximal set $Z_{i}\subset[n]$ such that there exists $x_{i}\in[m]^{Z_{i}}$ with the following properties:
$\mu_{Z_{i}\to x_{i}}(\mathcal{G}_{i-1}(Z_{i}\to x_{i}))\geqslant\tau^{|Z_{i}|}\mu(\mathcal{G}_{i-1}).$
If such $Z_{i}$ does not exist, set $l=i$, $Z_{i}=\varnothing$, $\mathcal{F}_{l}=\mathcal{G}_{l-1}$. Otherwise, if $|Z_{i}|>q$ or $\mu_{Z_{i}\to x_{i}}(\mathcal{G}_{i-1}(Z_{i}\to x_{i}))\leqslant\tau^{-q}$, then stop, put l=i1l=i-1 and set $\mathcal{R}=\mathcal{G}_{i-1}$. Finally, if $|Z_{i}|\leqslant q$ and $\mu_{Z_{i}\to x_{i}}(\mathcal{G}_{i-1}(Z_{i}\to x_{i}))\geqslant\tau^{-q}$, then set $\mathcal{F}_{i}=\mathcal{G}_{i-1}[Z_{i}\to x_{i}]$, $\mathcal{G}_{i}=\mathcal{G}_{i-1}\setminus\mathcal{F}_{i}$ and move to the step $i+1$.
In the end of procedure, we have either $\mu(\mathcal{R})\leqslant\tau^{-|Z_{l+1}|}\mu_{Z_{l+1}\to r_{l+1}}(\mathcal{G}_{l}(Z_{l+1}\to x_{l+1}))$ for some $|Z_{l+1}|>q$ or $\mu(\mathcal{R})\leqslant\tau^{-q}$. In both cases, we have $\mu(\mathcal{R})\leqslant\tau^{-q}$. By Lemma 13, families $\mathcal{F}_{i}$ are $\tau$-homogeneous, and the proof is complete. ∎
The core idea of our proof is to show that an extremal family $\mathcal{F}\subset[m]^{n}$ avoiding an agreement of size exactly t1t-1 is $t$-agreeing. At the first stage, it is enough to show that restrictions $(Z_{i},x_{i}),i=1,\ldots,l$, of the spread approximation are $t$-agreeing, so $\mathcal{F}$ is $t$-agreeing up to small remainder. Arguing indirectly, suppose that there are two restrictions $(Z_{1},x_{1})$ and $(Z_{2},x_{2})$ in $\{(Z_{i},x_{i})\}_{i=1}^{l}$ such that $\operatorname{agr}_{Z_{1}\cap Z_{2}}(x_{1},x_{2})<t$. Then, we will find a set $H\subset[n]$ of size $t-1-\operatorname{agr}_{Z_{1}\cap Z_{2}}(x_{1},x_{2})$ and a vector $y\in[m]^{H}$ such that such that $\operatorname{agr}_{(Z_{1}\cap Z_{2})\cup H}((x_{1},y),(x_{2},y))=t-1$, and some families $\mathcal{F}_{1}^{\prime}\subset\mathcal{F}(Z_{1}\cup H\to(x_{1},y))$, $\mathcal{F}_{2}\subset\mathcal{F}(Z_{2}\cup H\to(x_{2},y))$ are $\tau^{\prime}$-homogeneous and have a moderately large measure, see Lemma 29 below. Then, the measure boosting argument of the previous section guarantees that they cross-contain a disagreement, a contradiction.
Thus, we require some tools to work with restrictions of $\tau$-homogeneous families. In particular, we need the following lemma.
Lemma 15.
Let $\mathcal{F}\subset[m]^{n}$ be $\tau$-homogenous with respect to a $b$-balanced product measure $\nu$. Let $R\subset[n]$. Let $X\subset[m]^{R}$ be a set such that $X=\prod_{i\in R}X_{i}$ for some $X_{i}\subset[m]$, $\sum_{i\in R}|X_{i}|\tau b<m$. Define
$\mathcal{F}^{\prime}=\{y\in\mathcal{F}|\forall x\in X\,\operatorname{agr}(x,y|_{R})=0\}.$
Then, $\nu(\mathcal{F}^{\prime})\geqslant(1-\frac{\sum_{i\in R}|X_{i}|\tau b}{m})\nu(\mathcal{F})$. Moreover, $\mathcal{F}^{\prime}$ is $\tau^{\prime}$-homogenous for
$\tau^{\prime}=\tau\cdot\left(1-\frac{\sum_{i\in R}|X_{i}|\tau b}{m}\right)^{-1}.$
Proof.
For any $i\in R,x\in X_{i}$ we have
$\nu(\mathcal{F}[i\rightarrow x])=\nu_{i}(x)\nu_{i\to x}(\mathcal{F}(i\rightarrow x))\leqslant\nu_{i}(x)\tau\nu(\mathcal{F}),$
where the last inequality holds by the $\tau$-homogeneity of $\mathcal{F}$. Since $\nu$ is $b$-balanced, we have $\nu_{i}(x)\leqslant\frac{b}{m}$. It implies
$\nu(\mathcal{F}^{\prime})$ $=\nu(\mathcal{F}\setminus\bigcup_{i\in R}\bigcup_{x\in X_{i}}\mathcal{F}[i\rightarrow x])\geqslant\nu(\mathcal{F})-\sum_{i\in R}\sum_{x\in X_{i}}\nu(\mathcal{F}[i\rightarrow x])$
$\geqslant\left(1-\frac{\sum_{i\in R}|X_{i}|\tau b}{m}\right)\cdot\nu(\mathcal{F}),$
and the first statement of the lemma holds. Let us check the second statement on $\tau^{\prime}$-homogeneity of $\mathcal{F}^{\prime}$. Consider arbitrary $Z\subset[n]$ and $r\in[m]^{Z}$. We have
$\nu_{Z\rightarrow r}(\mathcal{F}^{\prime}(Z\rightarrow r))$ $\leqslant\nu_{Z\rightarrow r}(\mathcal{F}(Z\rightarrow r))\leqslant\tau^{|Z|}\nu(\mathcal{F})\leqslant\frac{\tau^{|Z|}}{\left(1-\frac{\sum_{i\in R}|X_{i}|\tau b}{m}\right)}\nu(\mathcal{F}^{\prime})$
$\leqslant\left(\frac{\tau}{1-\frac{1}{m}\sum_{i\in R}|X_{i}|\tau b}\right)^{|Z|}\nu(\mathcal{F}^{\prime})$
and, therefore, $\mathcal{F}^{\prime}$ is $\tau^{\prime}$-homogenous. ∎
To find certain restrictions, we widely use a simple averaging argument, which we state as a separate claim.
Claim 16.
Let $\mathcal{F}$ be a subset of $m]^{n}$ and let $\nu(\cdot)$ be a product measure on $m]^{n}$. Then, for any $H\subset[n]$ there exists $x\in[m]^{H}$ such that
$\nu_{H\to x}(\mathcal{F}(H\to x))\geqslant\nu(\mathcal{F}).$
Proof.
Consider a random vector $\mathbf{x}$ distributed on $m]^{H}$ according to the measure $\prod_{i\in S}\nu_{i}$. Then, we have
$\mathbb{E}\nu_{H\to\mathbf{x}}(\mathcal{F}(H\to\mathbf{x}))=\sum_{x\in[m]^{H}}\prod_{i\in H}\nu_{i}(x_{i})\cdot\frac{\nu(\mathcal{F}[H\to x])}{\prod_{i\in H}\nu_{i}(x_{i})}=\sum_{x\in[m]^{H}}\nu(\mathcal{F}[H\to x])=\nu(\mathcal{F}),$ (3)
and the claim follows. ∎
Another helpful argument that is based on similar ideas is the following lemma, which is applied to a pair of $\tau$-homogenous families simultaneously and allows to find a common restriction such that both restricted families have large measure.
Lemma 17.
Let $\mathcal{F}\subset[m]^{n}$ be a $\tau$-homogenous family with respect to a product measure $\nu$. Let $H\subset[n]$ and let $\mathbf{x}\sim\nu_{H}$ be a random element in $m]^{H}$. Let $p\in(0,1)$. If $\tau<\left(\frac{1+p}{2p}\right)^{\frac{1}{|H|}}$, then
$\mathbb{P}\left(\nu_{H\rightarrow\mathbf{x}}(\mathcal{F}(H\rightarrow\mathbf{x}))\geqslant\frac{1}{2}\nu(\mathcal{F})\right)>p.$
Proof.
Let $q=\mathbb{P}\left(\nu_{H\rightarrow\mathbf{x}}(\mathcal{F}(H\rightarrow\mathbf{x}))\geqslant\frac{1}{2}\nu(\mathcal{F})\right)$. We count $\mathbb{E}\nu_{H\rightarrow\mathbf{x}}(\mathcal{F}(H\rightarrow\mathbf{x}))$ in two ways. On the one hand, since $\nu_{H\rightarrow x}(\mathcal{F}(H\rightarrow x))\leqslant\tau^{|H|}\nu(\mathcal{F})$ for all $x$, we have
$\mathbb{E}\nu_{H\rightarrow\mathbf{x}}(\mathcal{F}(H\rightarrow\mathbf{x}))$ $=\mathbb{E}\nu_{H\rightarrow\mathbf{x}}(\mathcal{F}(H\rightarrow\mathbf{x}))\cdot\mathds{1}\left\{\nu_{H\rightarrow\mathbf{x}}(\mathcal{F}(H\rightarrow\mathbf{x}))\geqslant\frac{1}{2}\nu(\mathcal{F})\right\}$
$\quad+\mathbb{E}\nu_{H\rightarrow\mathbf{x}}(\mathcal{F}(H\rightarrow\mathbf{x}))\mathds{1}\left\{\nu_{H\rightarrow x}(\mathcal{F}(H\rightarrow x))<\frac{1}{2}\nu(\mathcal{F})\right\}$
$\leqslant q\tau^{|H|}\nu(\mathcal{F})+(1-q)\frac{1}{2}\nu(\mathcal{F}).$
On the other hand, we have $\mathbb{E}\nu_{H\rightarrow\mathbf{x}}(\mathcal{F}(H\rightarrow\mathbf{x}))=\nu(\mathcal{F})$ due to (3). Thus, we get $1\leqslant q\tau^{|H|}+\frac{1}{2}(1-q)$. It implies
$q\geqslant\frac{1}{2\tau^{|H|}-1}.$
Substituting $\tau<\left(\frac{1+p}{2p}\right)^{\frac{1}{|H|}}$, we get the desired bound $q>p$. ∎
Assume for a while, that pairs $(Z_{i},x_{i})$ obtained by the spread approximation lemma 14 are $t$-agreeing, i.e. for any $(Z_{i},x_{i}),(Z_{j},x_{j})$ from $\{(Z_{i},x_{i})\}_{i=1}^{l}$, we have $\operatorname{agr}_{{Z_{i}\cap Z_{j}}}(x_{i},x_{j})\geqslant t$. In the next section, we discuss a number of arguments due to papers [52] and [50], which help to deal with this and weaker properties of the spread approximations and to prove that there exists a pair $(T,x)$, $|T|=t$ and $x\in[m]^{T}$, such that for almost all codes $y\in\mathcal{F}$, we have $y|_{T}=x$.
3.3 Simplification arguments
Simplification arguments are one of the key components of the spread approximation technique, but they vary from problem to problem [52, 53, 55, 27, 50, 38]. All of them are designed for families of sets from an ambient family $\mathcal{A}$ that meet some conditions. First, we list them.
Say that a family of sets $\mathcal{F}$ is $r$-spread if for any $Z$ we have
$|\mathcal{F}(Z)|\leqslant r^{-|Z|}|\mathcal{F}|.$
Then, we require the following assumption, which is a stronger version of Assumption 1 [50].
Assumption 1.
A family $\mathcal{A}$ is $k$-uniform. Also, for any $X$ that is a subset of some set from $\mathcal{A}$ the family $\mathcal{A}(X)$ is $r$-spread.
The second assumption is used when, for a $\tau$-homogenous family $\mathcal{F}$, one needs to find a lot of sets $H$ in the shadow of $\mathcal{A}$ such that the $\mathcal{F}(H)$ is $\tau^{\prime}$-homogeneous for some $\tau^{\prime}$.
Assumption 2.
For any set $S\in\partial_{\leqslant q}\mathcal{A}$, a family $\mathcal{F}\subset\mathcal{A}(S)$, and $h\leqslant t-1$, we have
$\frac{|\mathcal{F}|}{|\mathcal{A}(S)|}=\frac{1}{|\partial_{h}(\mathcal{A}(S))|}\sum_{H\in\partial_{h}(\mathcal{A}(S))}\frac{|\mathcal{F}(H)|}{|\mathcal{A}(H\cup S)|},$
or, equivalently, for a random set $\mathbf{H}$ uniformly distributed on $\partial_{h}(\mathcal{A}(S))$, we have $\mu(\mathcal{F})=\mathbb{E}\mu(\mathcal{F}(\mathbf{H}))$, where $\mu$ is the uniform measure.
We proceed with verifying that the necessary assumptions hold for the family $\mathcal{C}$ defined by (1) with proper choice of $r,t,q,k$.
Lemma 18.
Assumptions 1 and 2 hold for the family $\mathcal{C}$, any positive real $r\leqslant m,$ integers $t\leqslant n+1,q\geqslant t$, and $k=n$.
Proof.
We start with verifying Assumption 1. Obviously, $\mathcal{C}$ considered as a subset of $mn]\choose n$ is $n$-uniform. Let $X$ be a subset of some set from $\mathcal{C}$, $Y$ be a subset of some set from $\mathcal{C}(X)$, i.e., $X$ and $Y$ are restrictions $(Z,x)$, $(W,y)$ for some disjoint $Z,W\subset[n],x\in[m]^{Z},y\in[m]^{W}$. We should prove that
$|\mathcal{C}(X\cup Y)|\leqslant r^{-|Y|}|\mathcal{C}(X)|.$
It is equivalent to
$m^{n-|X|-|Y|}\leqslant r^{-|Y|}\cdot m^{n-|X|},$
which holds for all $r\leqslant m$.
Now we verify Assumption 2. Let $S=(Z,x)$ be a restriction, $|Z|\leqslant q$ and let $\mathcal{F}$ be a subfamily of $m]^{n}(Z\to x)$. The expression to verify can be rewritten as follows
$\frac{|\mathcal{F}|}{m^{n-|Z|}}=\frac{1}{{n-|Z|\choose h}m^{h}}\sum_{H\in\partial_{h}([m]^{n}(Z\to x))}\frac{|\mathcal{F}(H)|}{m^{n-h-|Z|}},$
which is equivalent to
$|\mathcal{F}|\cdot{n-|Z|\choose h}=\sum_{H\in\partial_{h}([m]^{n}(Z\to x))}|\mathcal{F}(H)|.$
This equality holds since it is a double-counting of $h$-subsets of sets in $\mathcal{F}$ with multiplicities. ∎
Later we will only use $r=m$ since we have verified the $m$-spreadness and $k=n$ since we have verified $n$-uniformity. Concerning the spread approximation discussed in the previous section, if restrictions
$\{(Z_{i},x^{i})\}_{i=1}^{l}\subset\partial_{\leqslant q}[m]^{n}$
are $t$-agreeing, then the corresponding sets $\widetilde{Z}_{i}\in\partial_{\leqslant q}\mathcal{C}$, $\widetilde{Z}_{i}\cap I_{j}=\{(j-1)m+(x^{j})_{i}\}$ for $j\in Z_{i}$ and $i=1,\ldots,l$, are $t$-intersecting. We say that a $t$-intersecting family $\widetilde{\mathcal{S}}=\{\widetilde{Z}_{i}\}$ is trivial, if there exists a set $\widetilde{T}$ of size $t$ such that $\widetilde{T}\subset\widetilde{Z}$ for all $\widetilde{Z}\in\widetilde{\mathcal{S}}$. It turns out that any non-trivial $t$-intersecting family of restrictions spans a much smaller number of sets than a single set of size $t$. This was shown in the paper [52] under Assumption 1. We use the following improved result due to [53].
Theorem 19 (Theorem 14 [53]).
Let $\varepsilon\in(0;1],n,r,q,t\geqslant 1$ be such that $\varepsilon r\geqslant 24q$. Let $\mathcal{A}\subset 2^{[n]}$ satisfy Assumption 1 with this $r$ and let $\mathcal{S}\subset\binom{[n]}{\leqslant q}$ be a non-trivial $t$-intersecting family. Then, there exists a set $T\in\partial_{t}\mathcal{A}$ such that
$|\mathcal{A}[\mathcal{S}]|\leqslant\varepsilon|\mathcal{A}(T)|.$
We refer to Theorem 19 as a simplification argument, since it implies that if $\{(Z_{i},x_{i})\}_{i=1}^{l}$ are restrictions of the spread approximation of a large enough family $\mathcal{F}\subset[m]^{n}$, $n\geqslant\operatorname{poly}(t)\log m$, which avoids an agreement of size t1t-1, then they can be simultaneously ‘‘simplified’’ to a restriction of size $t$. In particular, Theorem 19 yields the following corollary in our setting.
Corollary 20.
Let $\varepsilon\in(0;1]$, $m,n,q,t\geqslant 1$ be such that $\varepsilon m\geqslant 24q$. Let $\mathcal{S}=\{(Z_{i},x_{i})\}_{i=1}^{l}\subset\partial_{\leqslant q}[m]^{n}$ be a family of $t$-agreeing restrictions. Suppose that there is no restriction $(T,x)$, $|T|=t$ and $x\in[m]^{T}$, such that $T\subset Z_{i}$ and $x_{i}|_{T}=x$ for all $i=1,\ldots,l$. Then, we have
$|[m]^{n}[\mathcal{S}]|\leqslant\varepsilon m^{n-t}.$
In the regime $n\geqslant\operatorname{poly}(t)\log m$, application of Corollary 20 almost completes the proof. It implies that, for an extremal family $\mathcal{F}\subset[m]^{n}$ that does not contain a (t1)(t-1)-agreement, there is a restriction $(T,x)$, $|T|=t$, such that $\mathcal{F}$ is almost fully contained in $\mathcal{F}[T\to x]$. Applying Lemma 12, we will deduce that $\mathcal{F}\setminus\mathcal{F}[T\to x]$ is empty.
Next, we move on to the discussion of the regime $m\geqslant n^{Ct}$ and tools required for its analysis, which were developed in the paper of Kupavskii and the second author [50]. First, we give some prelimanries concerning sunflowers and $(\mathcal{S},s,t)$-systems. Following Definition 8 of [50], we will name a family $\mathcal{S}=\{(Z_{i},x_{i})\}_{i=1}^{l}$ obtained from the spread approximation in Lemma 14.
Definition 4.
A family $\mathcal{S}=\{(Z_{1},x_{1}),\ldots,(Z_{l},x_{l})\}$ where $Z_{i}\subset[n]$ and $x_{i}\in[m]^{Z_{i}}$ is a $\tau$-homogenous core of $\mathcal{F}\subset[m]^{n}$ if $\mathcal{F}$ admits a decomposition
$\mathcal{F}=\bigsqcup_{S\in\mathcal{S}}\mathcal{F}_{S}$
where for all $S=(Z,x)\in\mathcal{S}$ we have $\mathcal{F}_{S}=\mathcal{F}_{S}[Z\to x]$ and families $\mathcal{F}_{S}(Z\to x)$ are $\tau$-homogenous.
We will sometimes treat $x_{i}$ as elements of $\partial_{\leqslant n}\mathcal{C}\subset\partial_{\leqslant n}{mn\choose n}$.
To tackle the case $m\geqslant n^{Ct}$, we will adapt the technique of $(\mathcal{S},s,t)$-systems of [50] which was developed for systems without sunflowers. Hence, we need to introduce the definition of a sunflower.
Definition 5.
A system of sets $F_{1},F_{2},\ldots,F_{s}$ is said to form a sunflower if and only if for all $i,j\in[s],i\neq j$ we have
$F_{i}\cap F_{j}=\bigcap_{k\in[s]}F_{k}.$
The set $\bigcap_{k\in[s]}F_{k}$ is called core, and sets $F_{1},F_{2},\ldots,F_{s}$ are called petals. After that we introduce the definition of $(\mathcal{S},s,t)$-system.
Definition 6.
Let $\mathcal{A}$ be an arbitrary family, $\mathcal{S}\subset\bigcup_{p=t}^{q}\partial_{p}\mathcal{A}$ be a family that does not contain a sunflower with $s$ petals and core of size t1t-1. We say that a family $\mathcal{B}\subset\mathcal{A}$ is an $(\mathcal{S},s,t)$-system, if $\mathcal{B}$ can be decomposed into disjoint families $\mathcal{B}_{S}\vee\{S\},S\in\mathcal{S}$ and $\mathcal{B}_{S}\subset\mathcal{A}(S)$, such that for any $s$ sets $S_{1},S_{2},\ldots,S_{s}\in\mathcal{S}$ the following properties hold.
  • $S_{1},S_{2},\ldots,S_{s}$ do not form a sunflower with the core of size exactly t1t-1;
  • if $S_{1},S_{2},\ldots,S_{s}$ form a sunflower with core of size at most t2t-2, then for any $F_{i}\in\mathcal{B}_{S_{i}},i\in[s]$ we have $|\bigcap_{i=1}^{s}F_{i}|\leqslant t-|S_{1}\cap S_{2}|-2$.
It is easy to check that our problem corresponds to forbidding a sunflower with $2$ petals and the core of size t1t-1 in a subfamily of $\mathcal{C}$. Hence, we may apply the main results of [50] on $(\mathcal{S},s,t)$-systems:
  1. 1.
    the decomposition of a family avoiding (t1)(t-1)-intersection into $(\mathcal{S},2,t)$-system;
  2. 2.
    the simplification of $(\mathcal{S},s,t)$-system into a moderately small number of incomplete $t$-stars.
Lemma 21 (Lemma 56 of [50]).
Let $\mathcal{F}\subset[m]^{n}$ be a code avoiding (t1)(t-1)-agreement, $t\leqslant n+1,q\geqslant t,n\geqslant q+t-1$. Let $\mathcal{S}$ be a $\tau$-homogenous core of $\mathcal{F}$ such that for all $(Z,x)=S\in\mathcal{S}$ we have $|Z|\leqslant q$. Let $\{\mathcal{F}_{S}\}_{S\in\mathcal{S}}$ be the corresponding decomposition of $\mathcal{F}$. Assume that $m>\max\{2^{15}\lceil\log_{2}n\rceil,2q\}\cdot 2\alpha^{1-t}\tau^{t}$ for some $\alpha\in(0,(4n)^{-1}].$ Then for all $(Z,x)=S\in\mathcal{S}$ there exist families $\mathcal{U}_{S}\subset\mathcal{F}_{S}=\mathcal{F}_{S}[Z\to x]$ such that
  1. 1.
    $|\mathcal{U}_{S}|\geqslant(1-2\alpha n)|\mathcal{F}_{S}|$ for each $S\in\mathcal{S}$;
  2. 2.
    $\mathcal{S}$ avoids intersection t1t-1; if $S_{1}=(Z_{1},x_{1}),S_{2}=(Z_{2},x_{2})\in\mathcal{S}$ agree on at most t2t-2 coordinates, then for any $y_{1}\in\mathcal{U}_{S_{1}},y_{2}\in\mathcal{U}_{S_{2}}$, we have $\operatorname{agr}_{[n]\setminus(Z_{1}\cup Z_{2})}(y_{1},y_{2})\leqslant t-|S_{1}\cap S_{2}|-2$;
  3. 3.
    let $\tau^{\prime}\leqslant\tau$ be the minimal homogenity of $\mathcal{F}_{S}$, then for any $h<n$ we have
    $|\partial_{h}(\mathcal{U}_{S})|\geqslant\left(\frac{\tau^{\prime}}{1-2\alpha n}\right)^{-h}|\partial_{h}([m]^{n}(S))|.$
It is easy to check that $\mathcal{U}_{S},S\in\mathcal{S},$ form a $(\mathcal{S},2,t)$-system. Moreover, the extra property 3 that bounds the shadow of each $\mathcal{U}_{S}$ below allows to prove the following lemma.
Lemma 22 (Proposition 57 of [50]).
Let $\mathcal{U}\subset[m]^{n}$ be an $(\mathcal{S},2,t)$-system. Let $\{\mathcal{U}_{S}\}_{S\in\mathcal{S}}$ be the decomposition of $\mathcal{U}$ from the definition of an $(\mathcal{S},s,t)$-system and $\lambda$ be a positive real number such that
$|\partial_{t-1}\mathcal{U}_{S}|\geqslant\lambda|\partial_{t-1}[m]^{n}|$
for every $S\in\mathcal{S}$. Suppose that $m\geqslant 8q^{3}$. Then there exists a family $\hat{\mathcal{T}}\subset\partial_{t}[m]^{n}$, such that
$|\hat{\mathcal{T}}|\leqslant\frac{1}{\lambda}(1+2\ln(\lambda|\partial_{\leqslant q}[m]^{n}|))$
and
$\sum_{S\in\mathcal{S}\setminus\mathcal{S}[\hat{\mathcal{T}}]}|\mathcal{U}_{S}|\leqslant\frac{32q^{3}}{\lambda m}\ln(\lambda|\partial_{\leqslant q}[m]^{n}|)\cdot m^{n-t}.$
Using spread approximation Lemma 14, one can decompose a (t1)(t-1)-avoiding code $\mathcal{F}$ into homogeneous families $\mathcal{F}_{S},S\in\mathcal{S}$, apply Lemma 21 and then deduce that $\mathcal{F}$ can be decomposed into relatively small number of incomplete $t$-stars $\mathcal{F}_{T},T\in\widehat{\mathcal{T}}$ up to some small remainder. This is a quite rich structure which we further refine. One of the key components of this part of our analysis is the Kruskal–Katona theorem for codes.
Theorem 23 ([22, 59]).
Suppose $\mathcal{F}\subset[m]^{n}$ and $\delta\geqslant 0$ is defined by $|\mathcal{F}|=\delta m^{n}$. Then for all $1\leqslant l\leqslant n$ one has $|\partial_{l}\mathcal{F}|\geqslant\delta^{\frac{l}{n}}|\partial_{l}[m]^{n}|$.
However, instead of the above theorem we often use its contraposition.
Corollary 24.
Suppose $\mathcal{F}\subset[m]^{n}$ and for some integer $l<n$ and some $\delta\geqslant 0$ we have $|\partial_{l}\mathcal{F}|\leqslant\delta|\partial_{l}[m]^{n}|$. Then one has $|\mathcal{F}|\leqslant\delta^{\frac{n}{l}}m^{n}$.
Applying Corollary 24 together with some geometrical scaling decomposition, we show that an extremal (t1)(t-1)-avoiding code $\mathcal{F}$ is almost fully contained in some $t$-star. Then, we apply Corollary 24 again and show that $\mathcal{F}$ must be a $t$-star, completing the case $m\geqslant n^{Ct}$.
4 Cross-disagreement in homogeneous families
In this section we prove a sufficient condition of existing of a cross-disagreement in homogenous families. The key ingredient of the proof is the following lemma.
Lemma 25.
Let $m,s$ be integer numbers such that $s|m$ and $s\geqslant 4$. Let $\nu$ be an $s$-balanced product measure on $m]^{n}$ and $\mathcal{F}$ be a $\tau$-homogenous family with respect to $\nu$. Then there exists a balanced gluing $\pi\in\Pi_{m,m/s}^{\otimes n}$ such that
$\nu^{\pi}(\mathcal{F}^{\pi})\geqslant\nu^{1-c(\tau)}(\mathcal{F}),$
where
$c(\tau)=\frac{\ln\frac{6}{5}}{8\ln\left(\frac{2^{7}}{\ln 4}\right)+4\ln\tau}.$
We will apply this lemma iteratively to transform exponentially small measures into constant measures.
Proof.
By Corollary 8 there exists $\pi\in\Pi_{m,m/s}^{\otimes n}$ such that $\nu^{\pi}(\mathcal{F}^{\pi})\geqslant\frac{\nu^{2}(\mathcal{F})}{\operatorname{Stab}_{5/6}(\mathcal{F})}$. Thus, it is sufficient to get an upper bound on $\operatorname{Stab}_{5/6}(\mathcal{F})$. By Proposition 9, an upper bound on $\operatorname{Stab}_{\rho_{0}}(\mathcal{F})$ for any $\rho_{0}>0$ implies an upper bound on $\operatorname{Stab}_{5/6}(\mathcal{F})$.
[…]
$\operatorname{Stab}_{\rho_{0}}(f)=\langle f,T_{\rho_{0}}f\rangle\leqslant\|T_{\rho_{0}}f\|_{4}\|f\|_{4/3}\leqslant\|f\|_{2}\|f\|_{4/3}=\nu^{5/4}(\mathcal{F}).$
Let $t$ be a number such that $t=2^{d},d\in\mathbb{N}$ and
$\frac{\ln\left(\frac{2^{7}}{\ln 4}\right)+\frac{1}{2}\ln\tau}{\ln\frac{6}{5}}\leqslant t<2\frac{\ln\left(\frac{2^{7}}{\ln 4}\right)+\frac{1}{2}\ln\tau}{\ln\frac{6}{5}}.$
[…]
$\operatorname{Stab}_{5/6}(f)\leqslant\|f\|^{2(1-1/t)}_{2}\operatorname{Stab}_{(\frac{5}{6})^{t}}(f)^{1/t}\leqslant\nu^{1-\frac{1}{t}}(\mathcal{F})\nu^{\frac{5}{4t}}(\mathcal{F})=\nu^{1+\frac{1}{4t}}(\mathcal{F})\leqslant\nu^{1+c(\tau)}(\mathcal{F}).$
Finally, we combine this bound with the inequality $\nu^{\pi}(\mathcal{F}^{\pi})\geqslant\frac{\nu^{2}(\mathcal{F})}{\operatorname{Stab}_{5/6}(\mathcal{F})}$ and get the desired bound $\nu^{\pi}(\mathcal{F}^{\pi})\geqslant\nu^{1-c(\tau)}(\mathcal{F})$. ∎
Now we are ready to proof a measure boosting lemma. The lemma is a counterpart of Lemma $6.11$ from [42].
Lemma 26.
Let $b\geqslant 2$ be an integer and $\nu$ be a $b$-balanced product measure on $m]^{n}$. Let $\mathcal{F}\subset[m]^{n},\nu(\mathcal{F})=\alpha$. Let $c(\tau)$ be as in Lemma 25,
$c_{1}(\tau)=4+\frac{2\ln\ln 2}{\ln(1-c(\tau))}$
and
$c_{2}(\tau)=-\frac{2}{\ln(1-c(\tau))}.$
For any $\tau>1$ such that
$m>b^{c_{1}(\tau)}\left(\ln\frac{1}{\alpha}\right)^{c_{2}(\tau)\ln b},$
and $n>\frac{\ln\alpha^{-1}}{\ln\tau}$ there are $m^{\prime}\in\mathbb{N},\pi\in\Pi_{m,m^{\prime},b}^{\otimes n},R\subset[n]$ and $x\in[m^{\prime}]^{R}$ such that
$m^{\prime}\geqslant\frac{m}{b^{c_{1}(\tau)}\left(\ln\frac{1}{\alpha}\right)^{c_{2}(\tau)\ln b}},$
$|R|\leqslant\frac{\ln\alpha^{-1}}{\ln\tau}$, $\mathcal{F}^{\pi}(R\rightarrow x)$ is $\tau$-homogenous with respect to $\nu^{\pi}_{R\rightarrow x}$ and $\nu^{\pi}_{R\rightarrow x}(\mathcal{F}^{\pi}(R\rightarrow x))\geqslant\frac{1}{2}$.
The main idea of the proof it to apply iteratively Lemma 25.
Proof.
Let $m_{0}$ be such a number, that $m_{0}=b^{k}$ for some $k\in\mathbb{N}$ and $m/b<m_{0}\leqslant m$. Let $\pi_{0}\in\Pi_{m,m_{0},b}^{\otimes n}$ be an arbitrary $b$-balanced gluing, $\nu^{\prime}_{0}=\nu^{\pi_{0}}$ and $\mathcal{F}^{\prime}_{0}=\mathcal{F}^{\pi_{0}}$. Note, that $\nu^{\prime}_{0}$ is $b^{2}$-balanced, sunce $\nu$ is $b$-balanced and $\pi_{0}$ is $b$-balanced. Let $Z_{0}\subset[n]$ be a maximal set such that for some $r_{0}\in[m_{0}]^{Z_{0}}$ we have $(\nu^{\prime}_{0})_{Z_{0}\to r_{0}}(\mathcal{F}^{\prime}_{0}(Z_{0}\to r_{0}))\geqslant\tau^{|Z_{0}|}\nu^{\prime}_{0}(\mathcal{F}^{\prime}_{0})$. Let $\nu_{0}=(\nu^{\prime}_{0})_{Z_{0}\to r_{0}}$ and $\mathcal{F}_{0}=\mathcal{F}^{\prime}_{0}(Z_{0}\to r_{0})$. By Lemma 13, the family $\mathcal{F}_{0}$ is $\tau$-homogenous with respect to $\nu_{0}$.
For $i\geqslant 1$ we iteratively construct tuples $(m_{i},\pi_{i},\nu_{i},Z_{i},r_{i},\mathcal{F}_{i})$ such that $m_{i}=\frac{m_{0}}{b^{2i}}$, $\pi_{i}\in\Pi_{m,m_{i},b}^{\otimes n}$, $Z_{i}\subset[n]$, $r_{i}\in[m_{i}]^{Z_{i}}$, $\nu_{i}=\nu^{\pi_{i}}_{Z_{i}\rightarrow r_{i}}$, $\mathcal{F}_{i}=\mathcal{F}^{\pi_{i}}(Z_{i}\to r_{i})$,
$\nu_{i}(\mathcal{F}_{i})\geqslant\max(\tau^{|Z_{i}|}\nu(\mathcal{F}),\nu(\mathcal{F})^{\left(1-c(\tau)\right)^{i}})$
and $\mathcal{F}_{i}$ is $\tau$-homogenous with respect to $\nu_{i}$. Note that $(m_{0},\pi_{0},\nu_{0},Z_{0},r_{0},\mathcal{F}_{0})$ satisfies the conditions. We construct the tuples by the following iterative procedure.
  1. 1.
    If $\nu_{i-1}(\mathcal{F}_{i-1})\geqslant\frac{1}{2}$ or $m_{i-1}\leqslant b$, stop.
  2. 2.
    Let $m_{i}=\frac{m_{i-1}}{b^{2}}$. Note that $b^{2}|m_{i-1}$, since $m_{i-1}$ is an integer power of $b$ and $m_{i-1}>b$. Let $\pi^{\prime}_{i}\in\Pi_{m_{i-1},m_{i}}^{\otimes([n]\setminus Z_{i-1})}$ be a gluing obtained by Lemma 25 with $s=b^{2}$ and $\nu=\nu_{i-1}$. Define $\nu^{\prime}_{i}=\nu_{i-1}^{\pi^{\prime}_{i}}$ and $\mathcal{F}^{\prime}_{i}=\mathcal{F}_{i-1}^{\pi^{\prime}_{i}}$. By Lemma 25, we get $\nu^{\prime}_{i}(\mathcal{F}^{\prime}_{i})\geqslant\nu_{i-1}(\mathcal{F}_{i-1})^{1-c(\tau)}$.
  3. 3.
    Consider an arbitrary balanced gluing $\pi_{i}^{\prime\prime}\in\Pi_{m_{i-1},m_{i}}^{\otimes Z_{i-1}}$. Let $\pi_{i}=(\pi_{i}^{\prime}\otimes\pi_{i}^{\prime\prime})\circ\pi_{i-1}$, $r^{\prime\prime}_{i}=\pi_{i}^{\prime\prime}(r_{i-1})$, $\mathcal{F}^{\prime\prime}_{i}=\mathcal{F}^{\pi_{i}}(Z_{i-1}\to r^{\prime\prime}_{i})$. Since $\mathcal{F}^{\prime}_{i}\subset\mathcal{F}^{\prime\prime}_{i}$, we get $\nu^{\prime}_{i}(\mathcal{F}^{\prime\prime}_{i})\geqslant\nu_{i-1}(\mathcal{F}_{i-1})^{1-c(\tau)}$. Moreover, $\pi_{i}$ is $b$-balanced, because $\pi_{i-1}$ is $b$-balanced and $(\pi_{i}^{\prime}\otimes\pi_{i}^{\prime\prime})$ is $1$-balanced.
  4. 4.
    Let $Z^{\prime}_{i}\subset[n]\setminus Z_{i-1}$ be maximal such that for some $r^{\prime}_{i}\in[m_{i}]^{Z^{\prime}_{i}}$ we have
    $(\nu^{\prime}_{i})_{Z^{\prime}_{i}\to r_{i}^{\prime}}(\mathcal{F}^{\prime\prime}_{i}(Z^{\prime}_{i}\to r^{\prime}_{i}))\geqslant\tau^{|Z^{\prime}_{i}|}\nu^{\prime}_{i}(\mathcal{F}^{\prime\prime}_{i}).$ (4)
    Let $\nu_{i}=(\nu^{\prime}_{i})_{Z^{\prime}_{i}\to r^{\prime}_{i}}$ and $\mathcal{F}_{i}=\mathcal{F}^{\prime\prime}_{i}(Z^{\prime}_{i}\to r^{\prime}_{i})$. By Lemma 13, the family $\mathcal{F}_{i}$ is $\tau$-homogenous with respect to $\nu_{i}$. Finally, let $Z_{i}=Z_{i-1}\sqcup Z^{\prime}_{i}$ and $r_{i}=(r^{\prime}_{i},r^{\prime\prime}_{i})$.
By construction, we get $m_{i}=\frac{m_{0}}{b^{2i}}$, $\pi_{i}\in\Pi_{m,m_{i},b}^{\otimes n}$, $Z_{i}\subset[n]$, $r_{i}\in[m_{i}]^{Z_{i}}$, $\nu_{i}=\nu^{\pi_{i}}_{Z_{i}\rightarrow r_{i}}$, $\mathcal{F}_{i}=\mathcal{F}^{\pi_{i}}(Z_{i}\to r_{i})$. By induction,
$\nu_{i}(\mathcal{F}_{i})\geqslant\nu^{\prime}_{i}(\mathcal{F}^{\prime\prime}_{i})\geqslant\nu_{i-1}(\mathcal{F}_{i-1})^{1-c(\tau)}\geqslant\left(\nu(\mathcal{F})^{\left(1-c(\tau)\right)^{i-1}}\right)^{1-c(\tau)}=\nu(\mathcal{F})^{\left(1-c(\tau)\right)^{i}}$
and
$\nu_{i}(\mathcal{F}_{i})\geqslant\tau^{|Z^{\prime}_{i}|}\nu^{\prime}_{i}(\mathcal{F}^{\prime\prime}_{i})\geqslant\tau^{|Z^{\prime}_{i}|}\nu_{i-1}(\mathcal{F}_{i-1})\geqslant\tau^{|Z^{\prime}_{i}|}\tau^{|Z_{i-1}|}\nu(\mathcal{F})=\tau^{|Z_{i}|}\nu(\mathcal{F}).$
Let $I$ be the value of $i$ when the procedure terminates. Let $m^{\prime}=m_{I-1},\pi=\pi_{I-1},R=Z_{I-1},x=r_{I-1}$. If $I=1$, then $m_{0}\geqslant m/b^{2}>b$, so $\nu_{0}(\mathcal{F}_{0})\geqslant 1/2$. If $I\geqslant 2$, we will show that the procedure terminates by the condition $\nu_{I-1}(\mathcal{F}_{I-1})\geqslant 1/2$ as well. Since $\alpha^{\left(1-c(\tau)\right)^{I-2}}\leqslant\nu_{I-2}(\mathcal{F}_{I-2})<\frac{1}{2}$, we get $I\leqslant 2+\frac{\ln\ln 2}{\ln(1-c(\tau))}-\frac{\ln\ln\frac{1}{\alpha}}{\ln(1-c(\tau))}$. Therefore,
$m_{I-1}\geqslant\frac{m}{b^{2I-1}}=b\frac{m}{b^{2I}}\geqslant b\frac{m}{b^{c_{1}(\tau)}\left(\ln\frac{1}{\alpha}\right)^{c_{2}(\tau)\ln b}}>b.$
Thus, we indeed have
$\nu^{\pi}_{R\rightarrow x}(\mathcal{F}^{\pi}(R\rightarrow x))=\nu_{I-1}(\mathcal{F}_{I-1})\geqslant\frac{1}{2}.$
It remains to bound $|R|$. Definition (4) ensures that $\tau^{|R|}\alpha\leqslant\nu_{I-1}(\mathcal{F}_{I-1})\leqslant 1$, so $|R|\leqslant\frac{\ln\alpha^{-1}}{\ln\tau}$. ∎
We are ready to prove a sufficient condition on existing of cross-disagreement in homogenous families.
Lemma 27.
Let $\mathcal{F}_{1}\subset[m]^{[n]\setminus H_{1}},\mathcal{F}_{2}\subset[m]^{[n]\setminus H_{2}}$. Let $b\geqslant 2$ and $\nu$ be a $b$-balanced product measure on $m]^{n}$. Let $\alpha_{i}=\nu_{[n]\setminus H_{i}}(\mathcal{F}_{i})$. Let $\mathcal{F}_{2}$ be $\tau$-homogenous with respect to $\nu_{[n]\setminus H_{2}}$. Let
$C_{1}(b)=\frac{4}{\ln 2}b^{3c_{1}(2)+4},\qquad C_{2}(b)=2c_{2}(2)\ln b+1,$
where $c_{1}(\tau)$ and $c_{2}(\tau)$ are from Lemma 26. If
$m>\tau C_{1}(b)\left(\max\left(1,\ln(\alpha_{1}^{-1})\right)\max\left(1,\ln(2\alpha_{2}^{-1})\right)\right)^{C_{2}(b)}$
and $n>\frac{\ln(\alpha_{1}^{-1})+\ln(2\alpha_{2}^{-1})}{\ln 2}+|H_{1}|+|H_{2}|$, then there exist $x_{i}\in\mathcal{F}_{i}$ such that $\operatorname{agr}_{[n]\setminus(H_{1}\cup H_{2})}\left(x_{1},x_{2}\right)=0$.
The formal proof of Lemma 27 is quite technical; therefore, we give its sketch first. We make several modifications of families $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$ to get families with much bigger measure on the same groud set $m^{\prime}]^{n^{\prime}}$ with $m^{\prime}\ll m$, such that if the modified families contain cross-disagreement than the initial families also contain cross-disagreement. Then we use Lemma 11 to find cross-disagreement in the modified families. The proof is divided into the following steps.
Step $1$. We use Lemma 26 to get a restriction $R\rightarrow x$ and a gluing $\pi$ such that $\mathcal{F}^{\prime}_{1}=\mathcal{F}_{1}^{\pi}(R\rightarrow x)$ has $\nu^{\pi}_{R\rightarrow x}$-measure at least $\frac{1}{2}$.
Step $2$. We erase from $\mathcal{F}_{2}$ sets, that agree after the gluing $\pi$ with $x$ on coordinates in $R$. Lemma 15 implies that the measure of $\mathcal{F}_{2}$ decreases insignificantly. We denote by $\mathcal{F}^{\prime}_{2}$ sets from $\mathcal{F}_{2}$, that were not erased in this step.
Step $3$. We transform the family $\mathcal{F}^{\prime}_{2}\subset[m]^{[n]\setminus H_{2}}$ into a family $\mathcal{F}^{\prime\prime}_{2}\subset[m^{\prime}]^{[n]\setminus(H_{2}\cup R)}$ without decreasing measure. First, using Claim 16, we find a restriction on $R\setminus H_{2}$ that does not decrease measure. Second, we apply a gluing $\pi$ obtained in the first step to this restriction. Claim 5 ensures that the gluing does not decrease measure. Note, that in this step we lose homogenity of $\mathcal{F}_{2}$. However, Lemma 26 ensures homogenity of $\mathcal{F}^{\prime}_{1}$.
Step $4$. We repeat steps 11-33, switching families. That is, we boost the measure of $\mathcal{F}^{\prime\prime}_{2}$ and ensure that the measure of $\mathcal{F}^{\prime}_{1}$ decreases insignificantly. After this step, we have families $\mathcal{F}^{\prime\prime\prime}_{1}\subset[m^{\prime\prime}]^{[n]\setminus(H_{1}\cup R\cup R^{\prime})}$ and $\mathcal{F}^{\prime\prime\prime}_{2}\subset[m^{\prime\prime}]^{[n]\setminus(H_{2}\cup R\cup R^{\prime})}$ with constant measures such that the existence of cross-disagreement in $\mathcal{F}^{\prime\prime\prime}_{1}$ and $\mathcal{F}^{\prime\prime\prime}_{1}$ implies the existence of cross-disagreement in the initial families.
Step $5$. At this step, the families $\mathcal{F}^{\prime\prime\prime}_{1}$ and $\mathcal{F}^{\prime\prime\prime}_{2}$ are in different sets of coordinates. We use Claim 16 to obtain restrictions $\mathcal{F}^{\prime\prime\prime\prime}_{1}$ and $\mathcal{F}^{\prime\prime\prime\prime}_{2}$ on the same ground set without decreasing the measure.
Step $6$. We use Lemma 11 to find the cross-disagreement in $\mathcal{F}^{\prime\prime\prime\prime}_{1}$ and $\mathcal{F}^{\prime\prime\prime\prime}_{2}$.
Next, we provide a formal proof of Lemma 27.
Proof.
Let $\beta_{1}=\min(\alpha_{1},e^{-1}),\beta_{2}=\min(\frac{\alpha_{2}}{2},e^{-1})$. Then, we have $\nu_{[n]\setminus H_{1}}(\mathcal{F}_{1})\geqslant\beta_{1},\nu_{[n]\setminus H_{2}}(\mathcal{F}_{2})\geqslant 2\beta_{2},\ln(\beta_{i}^{-1})\geqslant 1$, and the requirement on $m$ is equivalent to
$m>\tau C_{1}(b)\left(\ln(\beta_{1}^{-1})\ln(\beta_{2}^{-1})\right)^{C_{2}(b)}.$
Step $1$. By Lemma 26 for $\mathcal{F}=\mathcal{F}_{1}$ and $\tau=2$, we obtain $m^{\prime}\in\mathbb{N},\pi_{-H_{1}}\in\Pi_{m,m^{\prime},b}^{\otimes[n]\setminus H_{1}},R\subset[n]\setminus H_{1}$ and $x\in[m^{\prime}]^{R}$ such that
$m^{\prime}\geqslant\frac{m}{b^{c_{1}(2)}\left(\ln\frac{1}{\beta_{1}}\right)^{c_{2}(2)\ln b}}\geqslant\frac{4}{\ln 2}\tau b^{2c_{1}(2)+4}\left(\ln\left(\beta_{2}^{-1}\right)\right)^{2c_{2}(2)\ln b+1}\left(\ln\beta_{1}^{-1}\right),$
$|R|\leqslant\frac{\ln\alpha_{1}^{-1}}{\ln 2}$, $\mathcal{F}_{1}^{\pi_{-H_{1}}}(R\rightarrow x)$ is $2$-homogenous with respect to $(\nu_{[n]\setminus H_{1}})^{\pi_{-H_{1}}}_{R\rightarrow x}$ and
$(\nu_{[n]\setminus H_{1}})^{\pi_{-H_{1}}}_{R\rightarrow x}(\mathcal{F}_{1}^{\pi_{-H_{1}}}(R\rightarrow x))\geqslant\frac{1}{2}.$
Let $\pi\in\Pi_{m,m^{\prime},b}^{\otimes n}$ be arbitrary such that $\pi_{[n]\setminus H_{1}}=\pi_{-H_{1}}$. Let $\mathcal{F}^{\prime}_{1}=\mathcal{F}_{1}^{\pi}(R\rightarrow x)$ and $\nu^{\prime}=\nu^{\pi}_{[n]\setminus R}$.
Step $2$. Since $\pi$ is $b$-balanced, we get $|\pi_{i}^{-1}(x_{i})|\leqslant\frac{bm}{m^{\prime}}$ for every $i\in[n]$. Let $\mathcal{F}^{\prime}_{2}$ be the faimily of sets that do not agree after the gluing $\pi$ with $x$ on coordinates in $R$, i. e.
$\mathcal{F}^{\prime}_{2}=\{y\in\mathcal{F}_{2}|\forall z\in\pi_{R\setminus H_{2}}^{-1}(x|_{R\setminus H_{2}})\,\operatorname{agr}_{R\setminus H_{2}}(z,y)=0\}.$
By Lemma 15, we get
$\nu_{[n]\setminus H_{2}}(\mathcal{F}^{\prime}_{2})\geqslant\left(1-\frac{|R|\frac{bm}{m^{\prime}}\tau b}{m}\right)\nu_{[n]\setminus H_{2}}(\mathcal{F}_{2})\geqslant\left(1-\frac{\tau b^{2}\ln\alpha_{1}^{-1}}{m^{\prime}\ln 2}\right)\alpha_{2}\geqslant\frac{1}{2}\alpha_{2}.$
Since $\beta_{2}\leqslant\frac{1}{2}\alpha_{2}$, we have $\nu_{[n]\setminus H_{2}}(\mathcal{F}^{\prime}_{2})\geqslant\beta_{2}$.
Step $3$. Take arbitrary $y\in[m^{\prime}]^{R\setminus H_{2}}$ such that $\nu^{\pi}_{[n]\setminus(R\cup H_{2})}({\mathcal{F}^{\prime}_{2}}^{\pi}(R\setminus H_{2}\rightarrow y))\geqslant\nu_{[n]\setminus H_{2}}^{\pi}({\mathcal{F}^{\prime}_{2}}^{\pi})$. Note, that such $y$ exists by Claim 16, and $\operatorname{agr}(x,y)=0$ because for all $y$ with $\operatorname{agr}(x,y)\neq 0$ we have ${\mathcal{F}^{\prime}_{2}}^{\pi}(R\setminus H_{2}\rightarrow y)=\emptyset$. Let $\mathcal{F}^{\prime\prime}_{2}={\mathcal{F}^{\prime}_{2}}^{\pi}(R\setminus H_{2}\rightarrow y)$. By Claim 5, we get
$\nu^{\prime}_{[n]\setminus(R\cup H_{2})}(\mathcal{F}^{\prime\prime}_{2})=\nu^{\pi}_{[n]\setminus(R\cup H_{2})}(\mathcal{F}^{\prime\prime}_{2})\geqslant\nu^{\pi}_{[n]\setminus H_{2}}({\mathcal{F}^{\prime}_{2}}^{\pi})\geqslant\nu_{[n]\setminus H_{2}}(\mathcal{F}^{\prime}_{2})\geqslant\frac{1}{2}\alpha_{2}\geqslant\beta_{2}.$
For any $x^{\prime}\in\mathcal{F}^{\prime}_{1},y^{\prime}\in\mathcal{F}^{\prime\prime}_{2}$ there exist $x\in\mathcal{F}_{1},y\in\mathcal{F}_{2}$ such that
$\operatorname{agr}_{[n]\setminus(H_{1}\cup H_{2})}\left(\pi(x),\pi(y)\right)=\operatorname{agr}_{[n]\setminus(H_{1}\cup H_{2}\cup R)}\left(x^{\prime},y^{\prime}\right)$
and, therefore,
$\operatorname{agr}_{[n]\setminus(H_{1}\cup H_{2})}\left(x,y\right)\leqslant\operatorname{agr}_{[n]\setminus(H_{1}\cup H_{2}\cup R)}\left(x^{\prime},y^{\prime}\right).$
Thus, it is sufficient to find $x^{\prime}\in\mathcal{F}^{\prime}_{1},y^{\prime}\in\mathcal{F}^{\prime\prime}_{2}$ with $\operatorname{agr}_{[n]\setminus(H_{1}\cup H_{2}\cup R)}\left(x^{\prime},y^{\prime}\right)=0$.
Step $4$. Since $\nu$ is $b$-balanced and $\pi\in\Pi_{m,m^{\prime},b}^{\otimes n}$, by Claim 4, $\nu^{\pi}$ is $b^{2}$-balanced and therefore $\nu^{\prime}$ is $b^{2}$-balanced. We apply Lemma 26 for $\mathcal{F}=\mathcal{F}^{\prime\prime}_{2}$, $b=b^{2}$, $\nu=\nu^{\prime}$ and $\tau=2$ and obtain $m^{\prime\prime}\in\mathbb{N},\pi^{\prime}_{-H_{2}}\in\Pi_{m^{\prime},m^{\prime\prime},b^{2}}^{\otimes[n]\setminus(R\cup H_{2})},R^{\prime}\subset[n]\setminus(R\cup H_{2})$ and $y^{\prime}\in[m^{\prime\prime}]^{R^{\prime}}$ such that
$m^{\prime\prime}\geqslant\frac{m^{\prime}}{b^{2c_{1}(2)}\left(\ln(\beta_{2}^{-1})\right)^{2c_{2}(2)\ln b}}\geqslant\frac{4}{\ln 2}b^{4}\ln(\beta_{2}^{-1}),$
$|R^{\prime}|<\frac{\ln\left(2\alpha_{2}^{-1}\right)}{\ln 2}$, and $(\nu^{\prime})^{\pi^{\prime}_{-H_{2}}}_{R^{\prime}\rightarrow y^{\prime}}({\mathcal{F}_{2}^{\prime\prime}}^{\pi^{\prime}_{-H_{2}}}(R^{\prime}\rightarrow y^{\prime}))\geqslant\frac{1}{2}$. Let $\pi^{\prime}\in\Pi_{m,m^{\prime},b}^{\otimes[n]\setminus R}$ be arbitrary such that $\pi^{\prime}_{[n]\setminus(R\cup H_{2})}=\pi^{\prime}_{-H_{2}}$. Let $\mathcal{F}^{\prime\prime\prime}_{2}={\mathcal{F}_{2}^{\prime\prime}}^{\pi^{\prime}}(R^{\prime}\rightarrow y^{\prime})$ and $\nu^{\prime\prime}=\nu^{\prime\pi^{\prime}}_{R^{\prime}\rightarrow y^{\prime}}=\nu^{\prime\pi^{\prime}}_{[n]\setminus(R\cup R^{\prime})}$.
Since $\pi^{\prime}$ is $b^{2}$-balanced, we get $|(\pi^{\prime}_{i})^{-1}(y^{\prime}_{i})|\leqslant\frac{b^{2}m^{\prime}}{m^{\prime\prime}}$ for all $i\in R^{\prime}$. Let $\mathcal{F}^{\prime\prime}_{1}$ be the family of sets from $\mathcal{F}^{\prime}_{1}$ that do not agree after the gluing $\pi^{\prime}$ with $y^{\prime}$ on coordinates in $R^{\prime}$, i. e.
$\mathcal{F}^{\prime\prime}_{1}=\{x\in\mathcal{F}^{\prime}_{1}|\forall z\in\left(\pi_{R^{\prime}\setminus H_{1}}^{\prime}\right)^{-1}(y^{\prime})\,\operatorname{agr}_{R^{\prime}\setminus H_{1}}\left(z,x\right)=0\}.$
By Lemma 15, we get
$\nu^{\prime}_{[n]\setminus(R\cup H_{1})}(\mathcal{F}^{\prime\prime}_{1})\geqslant\left(1-\frac{2|R^{\prime}|\frac{b^{2}m^{\prime}}{m^{\prime\prime}}b^{2}}{m^{\prime}}\right)\nu^{\prime}_{[n]\setminus(R\cup H_{1})}(\mathcal{F}^{\prime}_{1})\geqslant\left(1-\frac{2b^{4}\ln\left(\beta_{2}^{-1}\right)}{m^{\prime\prime}\ln 2}\right)\cdot\frac{1}{2}\geqslant\frac{1}{4}.$
Take arbitrary $x^{\prime}\in[m^{\prime\prime}]^{R^{\prime}\setminus H_{1}}$ such that $\nu^{\prime\pi^{\prime}}_{[n]\setminus(R\cup R^{\prime}\cup H_{1})}({\mathcal{F}^{\prime\prime}_{1}}^{\pi^{\prime}}(R^{\prime}\setminus H_{1}\rightarrow x^{\prime}))\geqslant\nu^{\prime\pi^{\prime}}_{[n]\setminus(R\cup H_{1})}({\mathcal{F}_{1}^{\prime\prime}}^{\pi^{\prime}})$. Note, that such $x^{\prime}$ exists by Claim 16 and $\operatorname{agr}(x^{\prime},y^{\prime})=0$, because for all $x^{\prime}$ with $\operatorname{agr}(x^{\prime},y^{\prime})\neq 0$ we have ${\mathcal{F}_{1}^{\prime\prime}}^{\pi^{\prime}}(R^{\prime}\setminus H_{1}\rightarrow x^{\prime})=\emptyset$. Let $\mathcal{F}^{\prime\prime\prime}_{1}={\mathcal{F}_{1}^{\prime\prime}}^{\pi^{\prime}}(R^{\prime}\setminus H_{1}\rightarrow x^{\prime})$. By Claim 5 we get
$\nu^{\prime\prime}_{[n]\setminus(R\cup R^{\prime}\cup H_{1})}(\mathcal{F}^{\prime\prime\prime}_{1})=\nu^{\prime\pi^{\prime}}_{[n]\setminus(R\cup R^{\prime}\cup H_{1})}(\mathcal{F}^{\prime\prime\prime}_{1})\geqslant\nu^{\prime\pi^{\prime}}_{[n]\setminus(R\cup H_{1})}({\mathcal{F}_{1}^{\prime\prime}}^{\pi^{\prime}})\geqslant\nu^{\prime}_{[n]\setminus(R\cup H_{1})}(\mathcal{F}^{\prime\prime}_{1})\geqslant\frac{1}{4}.$
For any $x^{\prime\prime}\in\mathcal{F}^{\prime\prime\prime}_{1},y^{\prime\prime}\in\mathcal{F}^{\prime\prime\prime}_{2}$, there exist $x^{\prime}\in\mathcal{F}^{\prime}_{1},y^{\prime}\in\mathcal{F}^{\prime\prime}_{2}$ such that
$\operatorname{agr}_{[n]\setminus(H_{1}\cup H_{2}\cup R)}\left(\pi^{\prime}(x^{\prime}),\pi^{\prime}(y^{\prime})\right)=\operatorname{agr}_{[n]\setminus(H_{1}\cup H_{2}\cup R\cup R^{\prime})}\left(x^{\prime\prime},y^{\prime\prime}\right),$
and, therefore,
$\operatorname{agr}_{[n]\setminus(H_{1}\cup H_{2}\cup R)}\left(x^{\prime},y^{\prime}\right)\leqslant\operatorname{agr}_{[n]\setminus(H_{1}\cup H_{2}\cup R\cup R^{\prime})}\left(x^{\prime},y^{\prime}\right).$
Thus, it is sufficient to find $x^{\prime\prime}\in\mathcal{F}^{\prime\prime\prime}_{1},y^{\prime\prime}\in\mathcal{F}^{\prime\prime\prime}_{2}$ with
$\operatorname{agr}_{[n]\setminus(H_{1}\cup H_{2}\cup R\cup R^{\prime})}\left(x^{\prime\prime},y^{\prime\prime}\right)=0.$
Step $5$. By Claim 16 there exist $z_{i}\in[m]^{(H_{1}\cup H_{2})\setminus(H_{i}\cup R\cup R^{\prime})}$ such that
$\nu^{\prime\prime}_{[n]\setminus(R\cup R^{\prime}\cup H_{1}\cup H_{2})}(\mathcal{F}^{\prime\prime\prime}_{i}((H_{1}\cup H_{2})\setminus(H_{i}\cup R\cup R^{\prime})\rightarrow z_{i}))\geqslant\nu^{\prime\prime}_{[n]\setminus(R\cup R^{\prime}\cup H_{i})}(\mathcal{F}^{\prime\prime\prime}_{i}).$
Let $\mathcal{F}^{\prime\prime\prime\prime}_{i}=\mathcal{F}^{\prime\prime\prime}_{i}((H_{1}\cup H_{2})\setminus(H_{i}\cup R\cup R^{\prime})\rightarrow z_{i})$. Note, that only at this step we get the same set of coordinates for both families.
Step $6$. Finally, we apply Lemma 11 for $\mathcal{F}^{\prime\prime\prime\prime}_{1}$ and $\mathcal{F}^{\prime\prime\prime\prime}_{2}$ to find $x^{\prime\prime}\in\mathcal{F}^{\prime\prime\prime\prime}_{1}$ and $y^{\prime\prime}\in\mathcal{F}^{\prime\prime\prime\prime}_{2}$ with $\operatorname{agr}(x^{\prime\prime},y^{\prime\prime})=0$. Since $\nu^{\prime}$ is $b^{2}$-balanced and $\pi^{\prime}\in\Pi_{m^{\prime},m^{\prime\prime},b^{2}}^{\otimes[n]\setminus R}$, by Claim 4 $\nu^{\prime\prime}$ is $b^{4}$-balanced. Thus, we get $\nu^{\prime\prime}_{i}(x)\leqslant\frac{b^{4}}{m^{\prime\prime}}\leqslant\frac{1}{3}$. Therefore, if $\mathcal{F}^{\prime\prime\prime\prime}_{1}$ and $\mathcal{F}^{\prime\prime\prime\prime}_{2}$ are cross-intersecting, we may apply Lemma 11 with $\lambda=\frac{1}{3}$ and get
$\frac{1}{8}\leqslant\nu^{\prime\prime}(\mathcal{F}^{\prime\prime\prime\prime}_{1})\nu^{\prime\prime}(\mathcal{F}^{\prime\prime\prime\prime}_{2})\leqslant\left(\frac{\lambda}{1-\lambda}\right)^{2}\left(1-\nu^{\prime\prime}(\mathcal{F}^{\prime\prime\prime\prime}_{1})\right)\left(1-\nu^{\prime\prime}(\mathcal{F}^{\prime\prime\prime\prime}_{2})\right)\leqslant\frac{3}{32}.$
However, $\frac{1}{8}>\frac{3}{32}$ and, therefore, there exist $x^{\prime\prime}\in\mathcal{F}^{\prime\prime\prime\prime}_{1},y^{\prime\prime}\in\mathcal{F}^{\prime\prime\prime\prime}_{2}$ with $\operatorname{agr}(x^{\prime\prime},y^{\prime\prime})=0$. Thus, there exist $(x_{R},x_{R^{\prime}},z_{1},x^{\prime\prime})\in\mathcal{F}_{1}$ and $(y_{R},y_{R^{\prime}},z_{2},y^{\prime\prime})\in\mathcal{F}_{2}$ such that $\pi_{R}(x_{R})=x,\pi_{R}(y_{R})=y,\pi^{\prime}_{R^{\prime}}(x_{R^{\prime}})=x^{\prime},\pi^{\prime}_{R^{\prime}}(y_{R^{\prime}})=y^{\prime}$ and $\operatorname{agr}(x,y)=\operatorname{agr}(x^{\prime},y^{\prime})=\operatorname{agr}(x^{\prime\prime},y^{\prime\prime})=0$. Therefore,
$\operatorname{agr}_{[n]\setminus(H_{1}\cup H_{2})}\left((x_{R},x_{R^{\prime}},z_{1},x^{\prime\prime}),(y_{R},y_{R^{\prime}},z_{2},y^{\prime\prime})\right)=0.$
5 Case $m\geqslant\operatorname{poly}(t)$, $n\geqslant\operatorname{poly}(t)\log m$
In this section, we prove the following theorem.
Theorem 28.
Let $t\geqslant 2$,
$n\geqslant t+\frac{6t^{2}\ln m}{\ln\left(\frac{9}{8}\right)}+2+\frac{6t\ln m+2\ln\left(\frac{32}{3}\right)}{\ln 2},$ (5)
$m\geqslant\frac{48}{\ln\left(\frac{9}{8}\right)}t^{3}\ln m+16t,$ (6)
$m\geqslant 2\left(\frac{3}{2}\right)^{2}C_{1}(2)\left(\ln\left(\frac{32}{3}m^{3t}\right)\right)^{2C_{2}(2)},$ (7)
where $C_{1}(b)$ and $C_{2}(b)$ are from Lemma 27 and
$m\geqslant 48\left(\frac{3}{\ln\frac{9}{8}}t^{2}\ln m+1\right).$ (8)
Let $\mathcal{F}\subset[m]^{n}$ be such family, that $\operatorname{agr}(x,y)\neq t-1$ for any $x,y\in\mathcal{F}$ and $|\mathcal{F}|\geqslant m^{n-t}$. Then there exist $Z\subset n$ and $x\in[m]^{Z}$ such that $|Z|=t$ and $\mathcal{F}=[m]^{n}[Z\rightarrow x]$.
The proof is divided into the following steps.
Step $1$ Using Lemma 14 we decompose $\mathcal{F}$ into homogenous families $\mathcal{F}[Z_{i}\rightarrow x_{i}]$, $i\in[l]$, with significant measures and a small residue $\mathcal{R}$.
Step $2$. Arguing indirectly, we prove that $\operatorname{agr}_{Z_{i}\cap Z_{j}}\left(x_{i},x_{j}\right)\geqslant t$ for all $i,j$. First, assuming the contrary, we use Lemma 15 to obtain subfamilies $\mathcal{F}^{\prime}_{i}\subset\mathcal{F}[Z_{i}\rightarrow x_{i}]$ and $\mathcal{F}^{\prime}_{j}\subset\mathcal{F}[Z_{j}\rightarrow x_{j}]$ such that $\operatorname{agr}_{Z_{i}\cup Z_{j}}\left(y_{i},y_{j}\right)=\operatorname{agr}_{Z_{i}\cap Z_{j}}\left(x_{i},x_{j}\right)$ for all $y_{i}\in\mathcal{F}^{\prime}_{i},y_{j}\in\mathcal{F}^{\prime}_{j}$. Second, we use Lemma 17 to find a common restriction $H\rightarrow y$ of families $\mathcal{F}^{\prime}_{i}$ and $\mathcal{F}^{\prime}_{j}$ on a set of coordinates $H$ with size $|H|=(t-1)-\operatorname{agr}_{Z_{i}\cap Z_{j}}\left(x_{i},x_{j}\right)$, which leaves the measures of the families significant. Finally, we use Lemma 27 to find a cross-disagreement in families $\mathcal{F}^{\prime}_{i}(Z_{i}\cup H\rightarrow(x_{i},y))$ and $\mathcal{F}^{\prime}_{j}(Z_{j}\cup H\rightarrow(x_{j},y))$. Since this cross-disagreement implies existence of codes, agreeing exactly in (t1)(t-1) coordinates, in $\mathcal{F}$, we conclude that $\operatorname{agr}_{Z_{i}\cap Z_{j}}\left(x_{i},x_{j}\right)\geqslant t$ for all $i,j$.
Step $3$. We use Corollary 20 to prove that there is $Z_{0}\subset[n]$ and $x_{0}\in[m]^{Z_{0}}$ such that $|Z_{0}|=t$ and for all $i\in[l]$ we have $Z_{0}\subset Z_{i}$ and $x_{0}=x_{i}|_{Z_{0}}$. That is, we prove that all the family $\mathcal{F}$, except the small residue $\mathcal{R}$, is contained in $m]^{n}[Z_{0}\rightarrow x_{0}]$ for some $Z_{0}$ with $|Z_{0}|=t$.
Step $4$. We prove that the residue $\mathcal{R}$ is empty, that is $\mathcal{F}\subset[m]^{n}[Z_{0}\rightarrow x_{0}]$. In this step we deal with huge family $\mathcal{F}[Z_{0}\rightarrow x_{0}]$ and small family $\mathcal{R}$, so we use Lemma 12 to find cross-disagreement in the unbalanced case.
The following lemma is a key ingredient of the second step of the proof of Theorem 28.
Lemma 29.
Let $\nu$ be a $b$-balanced product measure on $m]^{n}$. For $i\in\{1,2\}$ let $Z_{i}\subset[n],x_{i}\in[m]^{Z_{i}}$ and $\mathcal{F}_{i}\in[m]^{[n]\setminus Z_{i}}$ are $\tau$-homogenous with respect to $\nu_{Z_{i}\rightarrow x_{i}}$. Let $\nu_{Z_{i}\rightarrow x_{i}}(\mathcal{F}_{i})=\alpha_{i}$. Let $t,q$ be integer numbers such that $|Z_{i}|\leqslant q$ for $i\in\{1,2\}$ and $\operatorname{agr}_{Z_{1}\cap Z_{2}}\left(x_{1},x_{2}\right)<t$. If
$\tau\leqslant\left(1-\frac{2qb}{m}\right)\left(\frac{3}{2}\right)^{\frac{1}{t-1}}$
then there exist $H\subset[n]\setminus(Z_{1}\cup Z_{2})$ with $|H|=t-1-\operatorname{agr}_{Z_{1}\cap Z_{2}}\left(x_{1},x_{2}\right)$, $y\in[m]^{H}$ and $\mathcal{F}^{\prime}_{i}\subset\mathcal{F}_{i}(H\rightarrow y)$ such that
$\nu_{Z_{i}\sqcup H\rightarrow(x_{i},y)}\left(\mathcal{F}^{\prime}_{i}\right)\geqslant\frac{1}{2}\left(1-\frac{q\tau b}{m}\right)\alpha_{i},$
$\mathcal{F}^{\prime}_{i}$ are $\tau^{\prime}$-homogenous with respect to $\nu_{Z_{i}\sqcup H\rightarrow(x_{i},y)}$ with
$\tau^{\prime}=2\left(\frac{\tau}{\left(1-\frac{q\tau b}{m}\right)}\right)^{t},$
and for any $z_{i}\in\mathcal{F}^{\prime}_{i}$ we have $\operatorname{agr}_{Z_{1}\cup Z_{2}\setminus Z_{i}}\left(z_{i},x_{3-i}\right)=0$.
Note, that the last condition guarantees that for any $z_{1}\in\mathcal{F}^{\prime}_{1},z_{2}\in\mathcal{F}^{\prime}_{2}$ we have
$\begin{split}&\operatorname{agr}((x_{1},y,z_{1}),(x_{2},y,z_{2}))=\operatorname{agr}_{Z_{1}\cap Z_{2}}\left(x_{1},x_{2}\right)+|H|+\operatorname{agr}_{[n]\setminus(Z_{1}\cup Z_{2}\cup H)}\left(z_{1},z_{2}\right)=\\ &t-1+\operatorname{agr}_{[n]\setminus(Z_{1}\cup Z_{2}\cup H)}\left(z_{1},z_{2}\right).\end{split}$
Proof.
Let $h=t-1-\operatorname{agr}_{Z_{1}\cap Z_{2}}\left(x_{1},x_{2}\right)$. Let $H\in\binom{[n]\setminus(Z_{1}\cup Z_{2})}{h}$ be arbitrary.
Let $\mathcal{F}^{\prime\prime}_{i}=\{y\in\mathcal{F}_{i}|\operatorname{agr}_{Z_{1}\cup Z_{2}\setminus Z_{i}}\left(y,x_{3-i}\right)=0\}$. By Lemma 15
$\nu_{Z_{i}\rightarrow x_{i}}(\mathcal{F}^{\prime\prime}_{i})\geqslant\left(1-\frac{q\tau b}{m}\right)\alpha_{i}$
and $\mathcal{F}^{\prime\prime}_{i}$ is $\tau^{\prime\prime}$-homogenous with $\tau^{\prime\prime}=\left(\frac{\tau}{\left(1-\frac{q\tau b}{m}\right)}\right)$. Since $\tau<2$, we get
$\tau^{\prime\prime}<\left(\frac{\tau}{\left(1-\frac{2qb}{m}\right)}\right)\leqslant\left(\frac{3}{2}\right)^{\frac{1}{t-1}}$
Let $\textbf{y}\sim\nu_{H}$ be a random element of $m]^{H}$. By Lemma 17 for $p=\frac{1}{2}$ we have
$\mathbb{P}\left(\nu_{[n]\setminus(Z_{i}\cup H)}\left(\mathcal{F}^{\prime\prime}_{i}(H\rightarrow\textbf{y})\right)\geqslant\frac{1}{2}\nu_{Z_{i}\rightarrow x_{i}}(\mathcal{F}^{\prime\prime}_{i})\right)>\frac{1}{2}.$
Therefore, there exists $y\in[m]^{H}$ such that
$\nu_{[n]\setminus(Z_{i}\cup H)}\left(\mathcal{F}^{\prime\prime}_{i}(H\rightarrow y)\right)\geqslant\frac{1}{2}\nu_{Z_{i}\rightarrow x_{i}}(\mathcal{F}^{\prime\prime}_{i})\geqslant\frac{1}{2}\left(1-\frac{q\tau b}{m}\right)\alpha_{i}$
for both $i\in\{1,2\}$. Let $\mathcal{F}^{\prime}_{i}=\mathcal{F}^{\prime\prime}_{i}(H\rightarrow y).$
We need to check that $\mathcal{F}^{\prime}_{i}$ are $\tau^{\prime}$-homogenous with respect to $\nu_{[n]\setminus(Z_{i}\cup H)}$ with
$\tau^{\prime}=2\left(\tau^{\prime\prime}\right)^{t}=2\left(\frac{\tau}{\left(1-\frac{q\tau b}{m}\right)}\right)^{t}.$
Indeed, for any $H^{\prime}\subset[n]\setminus(Z_{i}\cup H)$ and $x^{\prime}\in[m]^{H^{\prime}}$, we get
$\begin{split}&\nu_{[n]\setminus(Z_{i}\cup H\cup H^{\prime})}\left(\mathcal{F}^{\prime}_{i}(H^{\prime}\rightarrow x^{\prime})\right)=\nu_{[n]\setminus(Z_{i}\cup H\cup H^{\prime})}\left(\mathcal{F}^{\prime\prime}_{i}(H^{\prime}\cup H\rightarrow(x^{\prime},y))\right)\leqslant\\ &\leqslant\tau^{\prime\prime|H^{\prime}|+|H|}\nu_{[n]\setminus Z_{i}}(\mathcal{F}^{\prime\prime}_{i})\leqslant 2\tau^{\prime\prime|H^{\prime}|+|H|}\nu_{[n]\setminus(Z_{i}\cup H)}(\mathcal{F}^{\prime}_{i})\leqslant\\ &\leqslant 2\tau^{\prime\prime|H^{\prime}|+t-1}\nu_{[n]\setminus(Z_{i}\cup H)}(\mathcal{F}^{\prime}_{i})\leqslant\left(2\tau^{\prime\prime t}\right)^{|H^{\prime}|}\nu_{[n]\setminus(Z_{i}\cup H)}(\mathcal{F}^{\prime}_{i}).\qed\end{split}$
The following lemma combines Lemmas 27 and 29 into the second step of the proof of Theorem 28.
Lemma 30.
Let $\mathcal{F}\subset[m]^{n}$ and $\nu$ be a $b$-balanced product measure on $m]^{n}$ for $b\geqslant 2$. For $i\in\{1,2\}$ let $Z_{i}\subset[n]$ and $x_{i}\in[m]^{Z_{i}}$. Let $t,q$ be integer numbers such that $|Z_{i}|\leqslant q$ for $i\in\{1,2\}$ and $\operatorname{agr}_{Z_{1}\cap Z_{2}}\left(x_{1},x_{2}\right)<t$. Let families $\mathcal{F}_{i}\subset\mathcal{F}[Z_{i}\rightarrow x_{i}]$ be such that $\mathcal{F}(Z_{i}\rightarrow x_{i})$ is $\tau$-homogenous with respect to $\nu_{Z_{i}\rightarrow x_{i}}$. Define $\alpha_{i}=\nu_{Z_{i}\rightarrow x_{i}}(\mathcal{F}_{i}(Z_{i}\rightarrow x_{i}))$. If
$\tau<\left(1-\frac{2qb}{m}\right)\left(\frac{3}{2}\right)^{\frac{1}{t-1}},$
$n>2q+t+\frac{\ln(\left(\frac{1}{2}\left(1-\frac{q\tau b}{m}\right)\alpha_{1}\right)^{-1})+\ln(2\left(\frac{1}{2}\left(1-\frac{q\tau b}{m}\right)\alpha_{2}\right)^{-1})}{\ln 2}$
and
$m>2\left(\frac{3}{2}\right)^{\frac{t}{t-1}}C_{1}(b)\left(\ln\left(2\left(\left(1-\frac{q\tau b}{m}\right)\alpha_{1}\right)^{-1}\right)\ln\left(4\left(\left(1-\frac{q\tau b}{m}\right)\alpha_{2}\right)^{-1}\right)\right)^{C_{2}(b)},$
where $C_{1}(b)$ and $C_{2}(b)$ are from Lemma 27, then there exist $x^{\prime}_{i}\in\mathcal{F}[Z_{i}\rightarrow x_{i}]$ such that $\operatorname{agr}(x^{\prime}_{1},x^{\prime}_{2})=t-1$.
Note, that $(Z_{1},x_{1})$ and $(Z_{2},x_{2})$ are not required to be distinct.
Proof.
Let $H\subset[n]\setminus(Z_{1}\cup Z_{2})$ and $y\in[m]^{H}$ be from Lemma 29. Define $\mathcal{F}^{\prime}_{i}=\mathcal{F}_{i}(H\cup Z_{i}\rightarrow(x_{i},y))$. Put $\alpha^{\prime}_{i}=\nu_{[n]\setminus(H\cup Z_{i})}(\mathcal{F}^{\prime}_{i})$ and
$\tau^{\prime}=2\left(\frac{\tau}{\left(1-\frac{q\tau b}{m}\right)}\right)^{t}<2\left(\frac{3}{2}\right)^{\frac{t}{t-1}}.$
By Lemma 29, families $\mathcal{F}^{\prime}_{i}$ are $\tau^{\prime}$-homogenous with respect to $\nu_{[n]\setminus(H\cup Z_{i})}$,
$\alpha_{i}^{\prime}\geqslant\frac{1}{2}\left(1-\frac{q\tau b}{m}\right)\alpha_{i}.$
Therefore, by Lemma 27, there exist $z_{i}\in\mathcal{F}^{\prime}_{i}$ such that $\operatorname{agr}_{[n]\setminus(H\cup Z_{1}\cup Z_{2})}\left(z_{1},z_{2}\right)=0$. Let $x_{i}^{\prime}=(x_{i},y,z_{i})$. By the last condition from Lemma 29 we get
$\operatorname{agr}_{Z_{2}\setminus Z_{1}}\left(z_{1},x_{2}\right)=\operatorname{agr}_{Z_{1}\setminus Z_{2}}\left(x_{1},z_{2}\right)=0.$
So we have $x_{i}^{\prime}\in\mathcal{F}[Z_{i}\rightarrow x_{i}]$ and
$\begin{split}&\operatorname{agr}(x^{\prime}_{1},x^{\prime}_{2})=|H|+\operatorname{agr}_{Z_{1}\cap Z_{2}}\left(x_{1},x_{2}\right)+\operatorname{agr}_{Z_{1}\setminus Z_{2}}\left(x_{1},z_{2}\right)+\operatorname{agr}_{Z_{2}\setminus Z_{1}}\left(z_{1},x_{2}\right)+\\ &+\operatorname{agr}_{[n]\setminus(H\cup Z_{1}\cup Z_{2})}\left(z_{1},z_{2}\right)=(t-1-\operatorname{agr}_{Z_{1}\cap Z_{2}}\left(x_{1},x_{2}\right))+\\ &+\operatorname{agr}_{Z_{1}\cap Z_{2}}\left(z_{1},z_{2}\right)+0+0+0=t-1.\end{split}$
The next lemma proves that if almost all family $\mathcal{F}$ is contained in $m]^{n}[Z\rightarrow x]$ for some $Z$ with size $t$, then all $\mathcal{F}$ is contained in $m]^{n}[Z\rightarrow x]$. We use it in the last step of the proof of Theorem 28.
Lemma 31.
Let $m\geqslant 8$. Let $\mathcal{F}\subset[m]^{n}$ be such a family that $\operatorname{agr}(x,y)\neq t-1$ for any $x,y\in\mathcal{F}$ and $|\mathcal{F}|\geqslant m^{n-t}$. Assume that there exists $Z\subset n$ and $x\in[m]^{Z}$ such that $|Z|=t$ and $|\mathcal{F}\setminus[m]^{n}[Z\rightarrow x]|\leqslant m^{n-3t}$. Then $\mathcal{F}=[m]^{n}[Z\rightarrow x]$.
Proof.
Let $\mathcal{F}_{1}=\mathcal{F}[Z\rightarrow x]$ and $\mathcal{F}_{2}=\mathcal{F}\setminus\mathcal{F}_{1}$. We need to prove $\mathcal{F}\subset[m]^{n}[Z\rightarrow x]$, that is, $\mathcal{F}_{2}=\emptyset$. Since $|\mathcal{F}|\geqslant m^{n-t}=|[m]^{n}[Z\rightarrow x]|$, it implies $\mathcal{F}=[m]^{n}[Z\rightarrow x]$. Let $\mathcal{\delta}=\mu(\mathcal{F}_{2}),$ where $\mu$ is the uniform measure of $m]^{n}$. The condition $|\mathcal{F}\setminus[m]^{n}[Z\rightarrow x]|\leqslant m^{n-3t}$ implies $\delta\leqslant m^{-3t}$.
If $n<3t$, the condition $|\mathcal{F}\setminus[m]^{n}[Z\rightarrow x]|\leqslant m^{n-3t}$ immediately implies $|\mathcal{F}_{2}|<1$ and, therefore, $\mathcal{F}_{2}=\emptyset$. Thus, we may assume $n\geqslant 3t>2t-1$, that is, $n]\setminus Z>t-1$. Let $Z^{\prime}\in{[n]\setminus Z\choose t-1}$ be an arbitrary subset of (t1)(t-1) elements of $n]\setminus Z$. By Claim 16 there exists $(y,y^{\prime})\in[m]^{Z}\times[m]^{Z^{\prime}}$ such that $\mu_{[n]\setminus(Z\cup Z^{\prime})}(\mathcal{F}_{2}(Z\cup Z^{\prime}\rightarrow(y,y^{\prime})))\geqslant\mu(\mathcal{F}_{2})=\delta$. Let $\mathcal{F}^{\prime}_{2}=(\mathcal{F}_{2}(Z\cup Z^{\prime}\rightarrow(y,y^{\prime}))$.
If $\mathcal{F}_{2}\neq\emptyset$, we have $\mathcal{F}_{2}[Z\cup Z^{\prime}\rightarrow(y,y^{\prime})]\neq\emptyset$ and therefore $y\neq x$. Thus, we have $0\leqslant\operatorname{agr}(x,y)\leqslant t-1$, that is $0\leqslant(t-1)-\operatorname{agr}(x,y)\leqslant t-1$. Since $m\geqslant 2$, it implies that there exists $x^{\prime}\in[m]^{Z^{\prime}}$ such that $\operatorname{agr}(x^{\prime},y^{\prime})=(t-1)-\operatorname{agr}(x,y)$ and, therefore, $\operatorname{agr}((x,x^{\prime}),(y,y^{\prime}))=t-1$. Fix arbitrary such $x^{\prime}$ and let $\mathcal{F}^{\prime}_{1}=\mathcal{F}_{1}(Z\cup Z^{\prime}\rightarrow(x,x^{\prime}))$. The assumption $\mathcal{F}\geqslant m^{n-t}$ implies $|[m]^{n}[Z\rightarrow x]\setminus\mathcal{F}[Z\rightarrow x]|=m^{n-t}-|\mathcal{F}_{1}|=m^{n-t}-|\mathcal{F}|+|\mathcal{F}_{2}|\leqslant|\mathcal{F}_{2}|=\delta m^{n}$. Since $m]^{n}[Z\cup Z^{\prime}\rightarrow(x,x^{\prime})]\setminus\mathcal{F}[Z\cup Z^{\prime}\rightarrow(x,x^{\prime})]\subset[m]^{n}[Z\rightarrow x]\setminus\mathcal{F}[Z\rightarrow x]$, we get $|\mathcal{F}^{\prime}_{1}|=m^{n-2t+1}-|[m]^{n}[Z\cup Z^{\prime}\rightarrow(x,x^{\prime})]\setminus\mathcal{F}[Z\cup Z^{\prime}\rightarrow(x,x^{\prime})]|\geqslant m^{n-2t+1}-\delta m^{n}$ and, therefore, $\mu_{[n]\setminus(Z\cup Z^{\prime})}(\mathcal{F}^{\prime}_{1})\geqslant 1-\delta m^{2t-1}.$
We claim that there is no $x^{\prime\prime}\in\mathcal{F}^{\prime}_{1},y^{\prime\prime}\in\mathcal{F}^{\prime}_{2}$ such that $\operatorname{agr}(x^{\prime\prime},y^{\prime\prime})=0$. Indeed, otherwise, we have $(x,x^{\prime},x^{\prime\prime})\in\mathcal{F},(y,y^{\prime},y^{\prime\prime})\in\mathcal{F}$ and $\operatorname{agr}((x,x^{\prime},x^{\prime\prime}),(y,y^{\prime},y^{\prime\prime}))=\operatorname{agr}((x,x^{\prime}),(y,y^{\prime}))+\operatorname{agr}(x^{\prime\prime},y^{\prime\prime})=t-1$. Thus, Lemma 12 implies that $\mu_{[n]\setminus(Z\cup Z^{\prime})}(\mathcal{F}^{\prime}_{1})+\mu_{[n]\setminus(Z\cup Z^{\prime})}(\mathcal{F}^{\prime}_{2})^{\log_{m}(2)}\leqslant 1.$ Substituting bounds $\mu_{[n]\setminus(Z\cup Z^{\prime})}(\mathcal{F}^{\prime}_{1})\geqslant 1-\delta m^{2t-1}$, $\mu_{[n]\setminus(Z\cup Z^{\prime})}(\mathcal{F}^{\prime})_{2}\geqslant\delta$ and $m\geqslant 8$, we get
$1-\delta m^{2t-1}+\delta^{1/3}\leqslant 1.$
Since $\delta\leqslant m^{-3t}$, we have
$1\geqslant 1-\frac{\delta\delta^{-2/3}}{m}+\delta^{1/3}=1+\left(1-\frac{1}{m}\right)\delta^{1/3},$
and, consequently, $\delta=0$. That is, we get $\mathcal{F}_{2}=\emptyset$, and, therefore, $\mathcal{F}=[m]^{n}[Z\rightarrow x]$. ∎
Proof of Theorem 28. Step $1$. Let $\{(Z_{i},x_{i})|i=1\ldots l\}$ be from Lemma 14 for
$\tau=\left(\frac{9}{8}\right)^{\frac{1}{t}}$
and
$q=\lfloor\log_{\tau}(m^{3t})\rfloor+1=\left\lfloor\frac{3}{\ln\frac{9}{8}}t^{2}\ln m\right\rfloor+1.$
Note, that we have $\mu_{Z_{i}\rightarrow x_{i}}(\mathcal{F}_{i}(Z_{i}\rightarrow x_{i}))\geqslant\tau^{-q}\geqslant\frac{m^{-3t}}{\tau}\geqslant\frac{1}{2}m^{-3t}$ for any $i\in[l]$ and $\mu(\mathcal{R})\leqslant\tau^{-q}<m^{-3t}$, that is, $|\mathcal{R}|<m^{n-3t}$.
Step $2$. Assume, for the sake of contradiction, that for some $i,j\in[l]$ we have $\operatorname{agr}_{Z_{i}\cap Z_{j}}\left(x_{i},x_{j}\right)<t$. We apply Lemma 30 for $(Z_{i},x_{i})$ and $(Z_{j},x_{j})$ with $\nu=\mu$, $b=2$, $\tau=\tau$ and $q=q$. Let us check its conditions. Since by requirement (6) we have
$\left(1-\frac{4q}{m}\right)^{t}\geqslant 1-\frac{4qt}{m}\geqslant 1-\frac{4t\left(\log_{\tau}(m^{3t})+1\right)}{m}=1-\frac{\frac{12}{\ln\left(\frac{9}{8}\right)}t^{3}\ln m+4t}{m}\geqslant\frac{3}{4},$
we get
$\tau=\left(\frac{3}{4}\right)^{\frac{1}{t}}\left(\frac{3}{2}\right)^{\frac{1}{t}}<\left(1-\frac{4q}{m}\right)\left(\frac{3}{2}\right)^{\frac{1}{t-1}}$
and therefore the assumption on $\tau$ from Lemma 30 is satisfied. Since in terms of Lemma 30 we have $(1-\frac{q\tau b}{m})\geqslant\frac{3}{4}$ and $\alpha_{i}\geqslant\frac{1}{2}m^{-3t}$, the right-hand side in the requirement on $m$ from this lemma does not exceed
$2\left(\frac{3}{2}\right)^{2}C_{1}(2)\left(\ln\left(\frac{32}{3}m^{3t}\right)\right)^{2C_{2}(2)}$
and, therefore, the condition (7) implies that the assumption on $m$ from Lemma 30 is satisfied. Finally, the right-hand side in the requirement on $n$ from Lemma 30 does not exceed
$2q+t+\frac{2\ln\left(\frac{32}{3}m^{3t}\right)}{\ln 2}\leqslant t+\frac{6t^{2}\ln m}{\ln\left(\frac{9}{8}\right)}+2+\frac{6t\ln m+2\ln\left(\frac{32}{3}\right)}{\ln 2}$
and, therefore, this assumption is satisfied by (5). Thus, we may apply Lemma 30 and get $x_{1},x_{2}\in\mathcal{F}$ with $\operatorname{agr}(x_{1},x_{2})=t-1$, a contradiction. Hence, for all $i,j\in[l]$ we have $\operatorname{agr}_{Z_{i}\cap Z_{j}}\left(x_{i},x_{j}\right)\geqslant t$.
Step $3$. Assume, again for the sake of contradiction, that there is no $Z\subset[n]$ and $x\in[m]^{Z}$ such that $|Z|=t$ and for all $i\in[l]$ we have $Z\subset Z_{i}$ and $x=x_{i}|_{Z}$. Requirement (8) implies that we may apply Corollary 20 for $\varepsilon=\frac{1}{2}$. Thus, we get
$\left|\bigcup_{i\in[l]}\mathcal{F}[Z_{i}\rightarrow x_{i}]\right|\leqslant\left|\bigcup_{i\in[l]}[m]^{n}[Z_{i}\rightarrow x_{i}]\right|\leqslant\frac{1}{2}m^{n-t}$
and, therefore, $\mathcal{F}\leqslant\frac{1}{2}m^{n-t}+m^{n-3t}$. Since it contradicts the assumption $|\mathcal{F}|\geqslant m^{n-t}$, we conclude, that there exist $Z\subset[n]$ and $x\in[m]^{Z}$ such that $|Z|=t$ and for all $i\in[l]$ we have $Z\subset Z_{i}$ and $x=x_{i}|_{Z}$. Therefore, we have $\mathcal{F}\subset\mathcal{F}[Z\rightarrow x]\cup\mathcal{R}$ and
$|\mathcal{F}\setminus[m]^{n}[Z\rightarrow x]|\leqslant|\mathcal{R}|\leqslant m^{n-3t}.$
Step $4$. Finally, we apply Lemma 31 and get $\mathcal{F}=[m]^{n}[Z\rightarrow x]$. ∎
Although explicitly stated conditions (6), (7), (8), (5) in Theorem 28 are convenient to follow the proof, when we apply the theorem, it may be difficult to check four lengthy requirements. Thus, we conclude this section with an easier to use qualitative corollary of Theorem 28.
Corollary 32.
There exist a polynomial $n_{0}(t)$ of degree $2$ and a polynomial $m_{0}(t)$ of degree at most $300$ such that the following holds. If $n\geqslant n_{0}(t)\ln(m)$, $m\geqslant m_{0}(t)$, and $\mathcal{F}\subset[m]^{n}$ is such a family that $\operatorname{agr}(x,y)\neq t-1$ for any $x,y\in\mathcal{F}$ and $|\mathcal{F}|\geqslant m^{n-t}$, then there exist $Z\subset n$ and $x\in[m]^{Z}$ such that $|Z|=t$ and $\mathcal{F}=[m]^{n}[Z\rightarrow x]$.
Proof.
We need to check that conditions of Corollary 32 imply the conditions of Theorem 28. The case $t=1$, that is, the case of intersecting family of codes, is considered in [30]. Thus, we may assume $t\geqslant 2$. Requirements (6), (7), (8) from Theorem 28 have a form $m\geqslant\text{poly}(t,\log m)$ and the polynomial from the requirement (7) has the highest degree $2C_{2}(2)<300$ in $t$. Thus, there exists a polynomial $m_{0}(t)$ of degree at most $300$ such that $m\geqslant m_{0}(t)$ implies requirements (6), (7), (8). Finally, for $m>e$ the condition (5) is a consequence of a condition $n>n_{0}(t)\ln(m)$, where
$n_{0}(t)=t+\frac{6t^{2}}{\ln\left(\frac{9}{8}\right)}+2+\frac{4t+2\ln\left(\frac{32}{3}\right)}{\ln 2}$
is a polynomial of degree $2$. ∎
6 Case $m\geqslant n^{\operatorname{poly}(t)},n\geqslant\operatorname{poly}(t)$
In this section, we once again treat $m]^{n}$ as a subset of $mn]\choose n$.
We prove the following theorem:
Theorem 33.
Let $1\leqslant t\leqslant n+1,n\geqslant 50t^{2},m\geqslant(2^{22}n)^{4t}$. Let $\mathcal{F}\subset[m]^{n}$ be such a family, that $\operatorname{agr}(x,y)\neq t-1$ for any $x,y\in\mathcal{F}$ and $|\mathcal{F}|\geqslant m^{n-t}$. Then $|\mathcal{F}|=m^{n-t}$, moreover, there exists $Z\subset n$ and $x\in[m]^{Z}$ such that $|Z|=t$ and $\mathcal{F}$ coincides with $m]^{n}[Z\rightarrow x]$.
To prove this theorem, we will reduce a maximal (t1)(t-1)-avoiding family $\mathcal{F}$ to an $(\mathcal{S},2,t)$-system. In Section 6.1, we will prove a lemma that uses this system to obtain some nice decomposition of $\mathcal{F}$ up to a small remainder. In Section 6.2, we will derive useful properties of this decomposition and special subfamilies of its parts. In Section 6.3, we will prove that, given such a decomposition, one may conclude that a large part of $\mathcal{F}$ is contained in $m]^{n}[T]$ for some $t$-element $T$. Lastly, in Section 6.4, we combine the results of previous subsections to prove the main theorem of this section.
6.1 Decomposition via an $(\mathcal{S},2,t)$-system
We will need the following corollary of Lemma 22. Later we will apply this corollary to an $(\mathcal{S},2,t)$-system $\mathcal{U}$ obtained using Lemma 21. The particular bounds on $n,m,q$ could be explained as follows. We start the proof of the main theorem with constructing a $\tau$-homogeneous core consisting of sets of uniformity not greater than $q$. Since not all sets of the original family contain an element of the core, we have to deal with a remainder $\mathcal{R}$. Our technique allows us to bound its measure by $\tau^{-q}$, and we aim to bound it as $|\mathcal{R}|\leqslant m^{n-2t}$, so we need $\tau^{q}>m^{2t}$.
[…]
Corollary 34.
Let $m,n,t,q$ be such integers that $t\geqslant 1,n\geqslant 50t^{2},m\geqslant(2^{22}n)^{4t},q=24t^{2}$. Let $\mathcal{U}\subset[m]^{n}$ be an $(\mathcal{S},2,t)$-system and $\{\mathcal{U}_{S}\}_{S\in\mathcal{S}}$ be the decomposition of $\mathcal{U}$ from the definition of an $(\mathcal{S},s,t)$-system satisfying (1 — 3) of Lemma 21 with $\alpha=(5n)^{-1},\tau=m^{1/12t}$. Then there exists a family $\widetilde{\mathcal{T}}\subset\partial_{t}[m]^{n}$, such that
$|\widetilde{\mathcal{T}}|\leqslant m^{1/3}\ln m.$ (9)
Moreover,
$\sum_{S\in\mathcal{S}\setminus\mathcal{S}[\widetilde{\mathcal{T}}]}|\mathcal{U}_{S}|\leqslant m^{n-t-1/3}.$ (10)
Proof.
We apply Lemma 22 directly with $\lambda=m^{-1/6}$. We start with verifying the assumptions. First we note that, in fact, $m\geqslant(2^{22}n)^{4t}\geqslant(2^{22}\cdot 50t^{2})^{4t}\geqslant 8\cdot 24^{3}\cdot t^{6}=8q^{3}$. Then we verify that the assumption of Lemma 22 holds for $\lambda$ chosen before. Put $\tau^{\prime}\leqslant\tau=m^{1/12t}$ to be the minimal homogenity of $\mathcal{F}_{S},S\in\mathcal{S}$. Since (3) of Lemma 21 is satisfied, we have
$|\partial_{t-1}(\mathcal{U}_{S})|\geqslant\left(\frac{\tau^{\prime}}{1-2\alpha n}\right)^{-(t-1)}|\partial_{t-1}([m]^{n}(S))|.$
For all $(Z,x)=S\in\mathcal{S}$ we have $|Z|\leqslant q$, and it is easy to see that $|\partial_{t-1}([m]^{n}(S))|\geqslant|\partial_{t-1}([m]^{n-q})|={n-q\choose t-1}\cdot m^{t-1}$, where the last equality is due to $n-q\geqslant t-1$. Also recall that $\tau^{\prime}\leqslant m^{1/12t}$, $\alpha=(5n)^{-1}$ and $|\partial_{t-1}([m]^{n})|={n\choose t-1}\cdot m^{t-1}$. Thus, we have
$|\partial_{t-1}(\mathcal{U}_{S})|\geqslant\left(\frac{3/5}{\tau^{\prime}}\right)^{t-1}\cdot{n-q\choose t-1}\cdot m^{t-1}\geqslant\left(\frac{3/5}{m^{1/12t}}\right)^{t-1}\cdot{n-q\choose t-1}\cdot m^{t-1}.$
It remains to show that
$\left(\frac{3/5}{m^{1/12t}}\right)^{t-1}\cdot{n-q\choose t-1}\geqslant m^{-1/6}{n\choose t-1}$
which is equivalent to
$m^{1/6}\geqslant\left(\frac{m^{1/12t}}{3/5}\right)^{t-1}\cdot\frac{n\cdot\ldots\cdot(n-t+2)}{(n-q)\cdot\ldots\cdot(n-q-t+2)}.$
For all $0\leqslant i\leqslant t-2$ we have (ni)/(niq)<2(n-i)/(n-i-q)<2, because $n\geqslant 50t^{2}\geqslant t-2+48t^{2}=t-2+2q$. Thus, it is enough to prove the following bound:
$m^{1/6}\geqslant\left(\frac{m^{1/12t}}{3/5}\right)^{t-1}\cdot 2^{t-1}.$
It is equivalent to
$m\geqslant m^{(t-1)/2t}\cdot\left(\frac{2^{6}\cdot 5^{6}}{3^{6}}\right)^{t-1},$
which is true since
$m\geqslant m^{1/2}\cdot m^{1/2}\geqslant m^{(t-1)/2t}\left(\frac{2^{6}\cdot 5^{6}}{3^{6}}\right)^{t-1}.$
Therefore, Lemma 22 is applicable. We obtain a set of restrictions $\widetilde{\mathcal{T}}$ such that for any $(Z,x)\in\widetilde{\mathcal{T}}$ we have $|Z|=t$. We have
$|\widetilde{\mathcal{T}}|\leqslant\frac{1}{\lambda}(1+2\ln(\lambda|\partial_{\leqslant q}[m]^{n}|))=\frac{1+2\ln{\left(m^{-1/6}\cdot\sum_{p=0}^{q}{n\choose p}m^{p}\right)}}{m^{-1/6}}\leqslant$
$\leqslant\frac{1+2\ln{\left(2m^{-1/6}\cdot n^{q}m^{q}\right)}}{m^{-1/6}}\leqslant\frac{1+2q(\ln m+\ln n)}{m^{-1/6}}\leqslant 5qm^{1/6}\ln m\leqslant m^{1/3}\ln m,$
where the last inequality holds because $m\geqslant(2^{22}n)^{4t}\geqslant(2^{22}\cdot 50t^{2})^{4t}\geqslant 5^{6}\cdot 24^{6}\cdot t^{12}=(5q)^{6}$. Also, Lemma 22 implies
$\sum_{S\in\mathcal{S}\setminus\mathcal{S}[\hat{\mathcal{T}}]}|\mathcal{U}_{S}|\leqslant\frac{32q^{3}}{\lambda m}\ln(\lambda|\partial_{\leqslant q}[m]^{n}|)\cdot m^{n-t}=\frac{32q^{3}}{m^{5/6}}\ln{\left(m^{-1/6}\cdot\sum_{p=0}^{q}{n\choose p}m^{p}\right)}\cdot m^{n-t}\leqslant$
$\leqslant\frac{128q^{4}\ln m}{m^{5/6}}\cdot m^{n-t}\leqslant m^{n-t-1/3}.$
The last inequality holds because
$m^{1/2}=m^{1/4}\cdot m^{1/4}\geqslant(2^{22}n)^{t}\cdot\frac{\ln m}{4}\geqslant 128q^{4}\ln m,$
where we used $m^{1/4}=e^{1/4\ln m}\geqslant 1+\frac{1}{4}\ln m$, $(2^{22}n)^{t}\geqslant 2^{9}(24t^{2})^{4}$ and
$(2^{22}n)^{t}\geqslant(2^{22}\cdot 50t^{2})^{t}\geqslant 2^{22t}t^{8}\geqslant 2^{9}\cdot 24^{4}\cdot t^{8}.$
6.2 Properties of incomplete $t$-stars
In the previous section, we constructed a set of restrictions $\widetilde{\mathcal{T}}$ which ‘represents’ well the original family. We now study the incomplete $t$-stars which are subsets of $\mathcal{F}[T]$ for $T\in\widetilde{\mathcal{T}}$. In this section we do not use properties of $\widetilde{\mathcal{T}}$ so we prove all the claims for arbitrary $T\in\partial_{t}[m]^{n}$, however, we will apply them only to $T\in\widetilde{\mathcal{T}}$.
The following definition is crucial in our proof.
Definition 7.
Let $\mathcal{F}$ be a subfamily of $m]^{n}$, $T\in\partial_{t}[m]^{n},\mathcal{F}_{T}\subset\mathcal{F}[T]$. We define $\mathcal{F}^{*}_{T}\subset\mathcal{F}(T)$ as
[…]
We will later see that $\mathcal{F}^{*}_{T}$ is usually of almost the same size as $\mathcal{F}_{T}(T)$, however its properties allow us to find pairs of elements of $\mathcal{F}$ with fixed agreement given their shadows with the same agreement. This idea will be illustrated in the proof of the following claim.
Recall that we treat $m]^{n}$ as a subset of $mn]\choose n$. Since the proof is based on techniques originally developed for sets we use the same notation. We now prove that if $\mathcal{F}$ is (t1)(t-1)-avoiding then $\partial_{t-1}\mathcal{F}^{*}_{T}$ are disjoint for distinct $T$.
Claim 35.
Let $m,n,t$ be arbitrary positive integers, $\mathcal{F}\subset[m]^{n}$ be (t1)(t-1)-avoiding. For distinct $T_{1},T_{2}\in\partial_{t}\mathcal{F}$ and each $F_{1}\in\mathcal{F}_{T_{1}}^{*},F_{2}\in\mathcal{F}^{*}_{T_{2}}$ we have $|F_{1}\cap F_{2}|<t-1$.
Proof.
Assume the contrary, there exist some distinct $T_{1},T_{2}\in\partial_{t}\mathcal{F}$ and $F_{1}\in\mathcal{F}^{*}_{T_{1}},F_{2}\in\mathcal{F}^{*}_{T_{2}}$ such that $|F_{1}\cap F_{2}|\geqslant t-1$. Put $T_{1}=(Z_{1},x_{1}),T_{2}=(Z_{2},x_{2}),F_{1}=([n]\setminus Z_{1},y_{1}),F_{2}=([n]\setminus Z_{2},y_{2})$.
Since $T_{1}$ and $T_{2}$ are distinct and $|Z_{1}|=|Z_{2}|=t$, for $p=\operatorname{agr}_{Z_{1}\cap Z_{2}}(x_{1},x_{2})$ we have $0\leqslant p\leqslant t-1$. Also, $\operatorname{agr}_{[n]\setminus(Z_{1}\cup Z_{2})}(y_{1},y_{2})\geqslant t-1$. Choose a set $Z\subset[n]\setminus(Z_{1}\cup Z_{2})$ of (t1p)(t-1-p) coordinates on which $y_{1}$ and $y_{2}$ agree, put $x=y_{1}|_{Z}=y_{2}|_{Z}$. Since $F_{1}\in\mathcal{F}^{*}_{T_{1}},|Z|\leqslant t-1$, we can choose $F_{1}^{\prime}\in\mathcal{F}_{T_{1}}(T_{1})$ such that it avoids $F_{2}\sqcup T_{2}$ outside of $Z$. By analogy, we choose $F_{2}^{\prime}\in\mathcal{F}_{T_{2}}(T_{2})$ such that it avoids $F_{1}^{\prime}\sqcup T_{1}$ outside of $Z$. Both $F_{1}^{\prime}\sqcup T_{1}$ and $F_{2}^{\prime}\sqcup T_{2}$ belong to $\mathcal{F}$, however
$|(F_{1}^{\prime}\sqcup T_{1})\cap(F_{2}^{\prime}\sqcup T_{2})|=|T_{1}\cap T_{2}|+|Z|=t-1,$
which leads to contradiction. ∎
Set $\delta_{0}=e^{-n/t^{2}}$. Recall that our goal is to show that a maximal (t1)(t-1)-avoiding finally should be a complete $t$-star. In order to show that, we will prove that at most one of the restrictions from $\widetilde{\mathcal{T}}$ covers a big part of the original family. Later, we will prove that in fact this restriction is the center of a $t$-star which coincides with the original family. We are ready to prove the following proposition.
Proposition 36.
Let $m,n,t$ be such positive integers that $n\geqslant 50t^{2}$. Let $\mathcal{F}\subset[m]^{n}$ be (t1)(t-1)-avoiding. Then there are no two distinct $T_{1},T_{2}\in\partial_{t}\mathcal{F}$ such that $\mu^{-T_{1}}(\mathcal{F}^{*}_{T_{1}})\geqslant\delta_{0}$ and $\mu^{-T_{2}}(\mathcal{F}^{*}_{T_{2}})\geqslant\delta_{0}$.
Proof.
Assume the contrary. Put $T_{1}=(Z_{1},x_{1}),T_{2}=(Z_{2},x_{2})$. By averaging (Claim 16), we find restrictions $y_{1}\in[m]^{Z_{2}\setminus Z_{1}}$ and $y_{2}\in[m]^{Z_{1}\setminus Z_{2}}$ such that
$\mu_{[n]\setminus Z_{1}\cup Z_{2}}(\mathcal{F}^{*}_{T_{1}}(Z_{2}\setminus Z_{1}\to y_{1}))\geqslant\delta_{0};\qquad\mu_{[n]\setminus Z_{1}\cup Z_{2}}(\mathcal{F}^{*}_{T_{2}}(Z_{1}\setminus Z_{2}\to y_{2}))\geqslant\delta_{0}.$
Put $\mathcal{G}_{1}=\mathcal{F}_{T_{1}}^{*}(Z_{2}\setminus Z_{1}\to y_{1}),\mathcal{G}_{2}=\mathcal{F}_{T_{2}}^{*}(Z_{1}\setminus Z_{2}\to y_{2})$. By Theorem 23, we have
$|\partial_{t-1}\mathcal{G}_{1}|\geqslant\delta_{0}^{\frac{t-1}{n-|Z_{2}\cup Z_{1}|}}\cdot|\partial_{t-1}m^{[n]\setminus Z_{1}\cup Z_{2}}|=$
$e^{-\frac{n(t-1)}{t^{2}(n-|Z_{2}\cup Z_{1}|)}}\cdot|\partial_{t-1}m^{[n]\setminus Z_{1}\cup Z_{2}}|>$
$\frac{1}{2}{n-|Z_{2}\cup Z_{1}|\choose t-1}m^{t-1}.$
The last inequality holds trivially for $t=1$. For $t\geqslant 2$ we have $n/t(n-2t)\leqslant 0.6<\ln(2)$ and therefore
$e^{-\frac{n(t-1)}{t^{2}(n-|Z_{2}\cup Z_{1}|)}}\geqslant e^{-\frac{n}{t(n-2t)}}>e^{-\ln 2}\geqslant\frac{1}{2}.$
Analogously, $|\partial_{t-1}\mathcal{G}_{2}|\geqslant\frac{1}{2}{n-|Z_{1}\cup Z_{2}|\choose t-1}m^{n-|Z_{1}\cup Z_{2}|}.$ Since $\partial_{t-1}\mathcal{G}_{1},\partial_{t-1}\mathcal{G}_{2}\subset\partial_{t-1}[m]^{n\setminus(Z_{1}\cup Z_{2})}$, the intersection of $\partial_{t-1}\mathcal{G}_{1}$ and $\partial_{t-1}\mathcal{G}_{2}$ is not empty which contradicts Claim 35. ∎
6.3 Geometrically scaling decomposition
In this section, we discuss a (t1)(t-1)-avoiding family $\mathcal{F}\subset[m]^{n}$ decomposed as $\mathcal{R}\sqcup\bigsqcup_{T\in\widetilde{\mathcal{T}}}\mathcal{F}_{T},$ where $\widetilde{\mathcal{T}}\subset\partial_{t}\mathcal{F}$, $\mathcal{F}_{T}\subset\mathcal{F}[T]$ for all $T\in\widetilde{\mathcal{T}}$ and $\mathcal{R}$ is a relatively small remainder. We will show that there is a restriction $T\in\widetilde{\mathcal{T}}$ such that almost all sets of $\mathcal{F}$ are contained in $\mathcal{F}_{T}\vee\{T\}.$ We begin with bounding the number of parts of the decomposition depending on their measure.
Claim 37.
Let $m,n,t$ be positive integers such that $n\geqslant 50t^{2}$. Let $\mathcal{F}=\bigsqcup_{T\in\widetilde{\mathcal{T}}}\mathcal{F}_{T}$ be a (t1)(t-1)-avoiding subfamily of $m]^{n}$, $\widetilde{\mathcal{T}}\in\partial_{t}\mathcal{F}$ and $\mathcal{F}_{T}\subset\mathcal{F}[T]$ for all $T\subset\widetilde{\mathcal{T}}$. Set $\delta_{0}=e^{-n/t^{2}}$, for $i\in\mathbb{N}$ put $\delta_{i}=\delta_{0}^{i+1}.$ Define
$\widetilde{\mathcal{T}}_{i}=\{T\in\widetilde{\mathcal{T}}\;:\;\mu^{-T}(\mathcal{F}^{*}_{T})\in(\delta_{i},\delta_{i-1}]\}.$
Then for any $\widetilde{\mathcal{T}}_{i}$ we have $|\widetilde{\mathcal{T}}_{i}|\leqslant 2\delta_{i}^{-(t-1)/(n-t)}$.
Proof.
Assume the contrary. From Theorem 23 for all $T\in\widetilde{\mathcal{T}}_{i}$ we have
$|\partial_{t-1}\mathcal{F}^{*}_{T}|\geqslant{n-t\choose t-1}m^{t-1}\cdot\delta_{i}^{\frac{t-1}{n-t}}.$
Claim 35 guarantees that the (t1)(t-1)-shadows of $\mathcal{F}^{*}_{T}$ are disjoint for different $T$. Thus the following series of inequality holds.
$\Bigg|\partial_{t-1}\bigcup_{T\in\widetilde{\mathcal{T}}}\mathcal{F}^{*}_{T}\Bigg|=\sum_{T\in\widetilde{T}}|\partial_{t-1}\mathcal{F}^{*}_{T}|\geqslant$
$\geqslant|\widetilde{\mathcal{T}_{i}}|\cdot{n-t\choose t-1}m^{t-1}\cdot\delta_{i}^{\frac{t-1}{n-t}}\geqslant 2\cdot{n-t\choose t-1}m^{t-1}>{n\choose t-1}m^{t-1}=|\partial_{t-1}[m]^{n}|.$
This leads us to contradiction. Note that $2{n-t\choose t-1}>{n\choose t-1}$ because
$\frac{{n\choose t-1}}{{n-t\choose t-1}}=\frac{n(n-1)\cdot\ldots\cdot(n-t+2)}{(n-t)\cdot\ldots\cdot(n-2t+2)}\leqslant\left(1+\frac{t}{n-2t+2}\right)^{t-1}<2,$
where the last inequality is due to $n\geqslant 50t^{2}\Longrightarrow n-2t+2\geqslant 2t^{2}$. ∎
Using Claim 37, we will show in two steps that there exists a set $T_{0}\in\widetilde{\mathcal{T}}$ such that $|\mathcal{F}\setminus\mathcal{F}_{T_{0}}^{*}|$ is relatively small. First, in Claim 38, we prove that the total size of families $\mathcal{F}_{T}^{*}$ except some $\mathcal{F}^{*}_{T_{0}}$ is at most $2\sqrt{\delta_{0}}m^{n-t}.$ Then, in Claim 39, we prove that $|\mathcal{F}\setminus\bigcup_{T\in\widetilde{\mathcal{T}}}\mathcal{F}^{*}_{T}|\leqslant m^{n-t-1/3}.$
Claim 38.
In terms of Claim 37 set $T_{0}\in\widetilde{\mathcal{T}}$ be such that $\mu^{-T_{0}}(\mathcal{F}_{T_{0}}^{*})$ is the biggest among $\mu^{-T}(\mathcal{F}_{T}^{*}),T\in\widetilde{\mathcal{T}}$. Then
$\sum_{T\in\widetilde{\mathcal{T}}}|\mathcal{F}_{T}^{*}|\leqslant|\mathcal{F}^{*}_{T_{0}}|+2m^{n-t}\sqrt{\delta_{0}}.$
Proof.
Fix any $T_{0}$ for which $\mu(\mathcal{F}_{T_{0}})$ attains the maximum. According to Claim 37, we have
$\sum_{T\in\widetilde{\mathcal{T}}}|\mathcal{F}_{T}^{*}|=|\mathcal{F}^{*}_{T_{0}}|+\sum_{i=1}^{\infty}\sum_{T\in\widetilde{\mathcal{T}}_{i}}|\mathcal{F}^{*}_{T_{i}}|\leqslant|\mathcal{F}^{*}_{T_{0}}|+2m^{n-t}\sum_{i=1}^{\infty}\delta_{i}^{-(t-1)/(n-t)}\delta_{i-1}=$
$|\mathcal{F}^{*}_{T_{0}}|+2m^{n-t}\sum_{i=1}^{\infty}\delta_{0}^{i-\frac{(i+1)(t-1)}{n-t}}.$
For any $i\in\mathbb{N}$, since nt>6(t1)n-t>6(t-1), the following inequalities hold:
$i-\frac{(i+1)(t-1)}{n-t}\geqslant i-\frac{i+1}{6}\geqslant\frac{2i}{3}$
and we can proceed estimating $\sum_{T\in\widetilde{\mathcal{T}}}|\mathcal{F}_{T}^{*}|$ as
$\sum_{T\in\widetilde{\mathcal{T}}}|\mathcal{F}_{T}^{*}|\leqslant|\mathcal{F}^{*}_{T_{0}}|+2m^{n-t}\sum_{i=1}^{\infty}\delta_{0}^{2i/3}=|\mathcal{F}^{*}_{T_{0}}|+2m^{n-t}\delta_{0}^{2/3}\cdot\frac{1}{1-\delta_{0}^{2/3}}\leqslant|\mathcal{F}^{*}_{T_{0}}|+2m^{n-t}\sqrt{\delta_{0}}$
The last inequality holds, because $\delta_{0}=e^{-n/t^{2}}<e^{-3}<0.1$. ∎
We will now bound $\sum_{T\in\widetilde{\mathcal{T}}}|\mathcal{F}_{T}(T)\setminus\mathcal{F}^{*}_{T}|$.
Claim 39.
Let $m,n,t$ be positive integers such that $m\geqslant(2^{22}n)^{4t},n\geqslant 50t^{2}$. Let $\mathcal{F}=\bigsqcup_{T\in\widetilde{\mathcal{T}}}\mathcal{F}_{T}$ be a (t1)(t-1)-avoiding subfamily of $m]^{n}$, $\widetilde{\mathcal{T}}\in\partial_{t}\mathcal{F}$, $\mathcal{F}_{T}\subset\mathcal{F}[T]$ for all $T\in\widetilde{\mathcal{T}}$ and $|\widetilde{\mathcal{T}}|\leqslant m^{1/3}\ln m$. Then, we have
$\sum_{T\in\widetilde{\mathcal{T}}}|\mathcal{F}_{T}(T)\setminus\mathcal{F}^{*}_{T}|\leqslant m^{n-t-1/3}.$
Proof.
Recall the definition of $\mathcal{F}_{T}^{*}$. The sets from $\mathcal{F}_{T}(T)\setminus\mathcal{F}_{T}^{*}$ can be described as follows: there is a restriction $(Z,x),Z\in[n],x\in[m]^{Z},|Z|\leqslant t-1$ such that the covering number of $\mathcal{F}_{T}(Z\to x)$ is less than $n+1$. It also means that the size of $\mathcal{F}_{T}(T)(Z\to x)$ is at most $(n+1)\cdot m^{n-t-|Z|-1}$, and we can write
$\sum_{T\in\widetilde{\mathcal{T}}}|\mathcal{F}_{T}(T)\setminus\mathcal{F}^{*}_{T}|\leqslant|\widetilde{\mathcal{T}}|\cdot\sum_{\begin{subarray}{c}Z\in[n],|Z|\leqslant t-1\\ x\in[m]^{Z}\end{subarray}}(n+1)\cdot m^{n-t-|Z|-1}\leqslant 2n|\widetilde{\mathcal{T}}|\cdot\sum_{Z\in[n],|Z|\leqslant t-1}m^{n-t-1}\leqslant$
$\leqslant 2n\cdot m^{1/3}\ln m\cdot\sum_{p=0}^{t-1}{n\choose p}\cdot m^{n-t-1}\leqslant 4n^{t}\ln m\cdot m^{n-t-2/3}\leqslant m^{n-t-1/3},$ (11)
where we used
$m\geqslant(2^{22}n)^{3t}\cdot(m^{1/12})^{3}\geqslant(2^{22}n)^{3t}\cdot\left(\frac{\ln m}{12}\right)^{3}\geqslant 4^{3}n^{3t}\cdot\ln^{3}m=(4n^{t}\ln m)^{3}$
in the last inequality. ∎
We are now ready to prove that for some $T_{0}$ all but a constant factor sets from $\mathcal{F}^{*}_{T_{0}}\vee\{T_{0}\}$ are contained in $\mathcal{F}$, i.e. $\mathcal{F}$ contains a large incomplete $t$-star.
Claim 40.
Let $m,n,t$ be positive integers such that $m\geqslant(2^{22}n)^{4t},n\geqslant 50t^{2}$. Suppose that a (t1)(t-1)-avoiding family $\mathcal{F}\subset[m]^{n},|\mathcal{F}|\geqslant m^{n-t}$, can be decomposed into families $\mathcal{R}\sqcup\bigsqcup_{T\in\widetilde{\mathcal{T}}}\mathcal{F}_{T}$ such that $\widetilde{\mathcal{T}}\subset\partial_{t}\mathcal{F}$, $\mathcal{F}_{T}\subset\mathcal{F}[T]$ for all $T\in\widetilde{\mathcal{T}}$ and $|\mathcal{R}|\leqslant 2m^{n-t-1/3},|\widetilde{\mathcal{T}}|\leqslant m^{1/3}\ln m$. Then, there are $T_{0}\in\widetilde{\mathcal{T}}$ and $\delta<1/e$ such that
$\big|\mathcal{F}^{*}_{T_{0}}\vee\{T_{0}\}\big|=(1-\delta)m^{n-t}.$
Proof.
The family $\mathcal{F}$ can be partitioned into
$\mathcal{F}=\mathcal{R}\sqcup\left(\bigsqcup_{T\in\widetilde{\mathcal{T}}}\mathcal{F}_{T}\setminus(\mathcal{F}^{*}_{T}\vee\{T\})\right)\sqcup\left(\bigsqcup_{T\in\widetilde{\mathcal{T}}}(\mathcal{F}^{*}_{T}\vee\{T\})\right).$ (12)
According to Claim 39, we have
$\Bigg|\mathcal{R}\sqcup\left(\bigsqcup_{T\in\widetilde{\mathcal{T}}}\mathcal{F}_{T}\setminus(\mathcal{F}^{*}_{T}\vee\{T\})\right)\Bigg|\leqslant|\mathcal{R}|+\Big|\bigcup_{T\in\widetilde{\mathcal{T}}}\mathcal{F}_{T}(T)\setminus\mathcal{F}^{*}_{T}\Big|\leqslant 3m^{n-t-1/3}.$
Claim 38 implies
$|\mathcal{F}|\leqslant 3m^{n-t-1/3}+|\mathcal{F}_{T_{0}}^{*}|+2m^{n-t}\sqrt{\delta_{0}}=|\mathcal{F}_{T_{0}}^{*}|+(2\sqrt{\delta_{0}}+3m^{-1/3})m^{n-t},$
where $\delta_{0}=e^{-n/t^{2}}$ and $T_{0}\in\widetilde{\mathcal{T}}$ is such that $\mu^{-T_{0}}(\mathcal{F}_{T_{0}}^{*})$ is the biggest among $\mu^{-T}(\mathcal{F}_{T}^{*}),T\in\widetilde{\mathcal{T}}$. Since $|\mathcal{F}|\geqslant m^{n-t}$, we have
$|\mathcal{F}^{*}_{T_{0}}|\geqslant(1-2\sqrt{\delta_{0}}-3m^{-1/3})m^{n-t}.$
Set $\delta>0$ such that $\big|\mathcal{F}^{*}_{T_{0}}\vee\{T_{0}\}\big|=(1-\delta)m^{n-t}$. According to $m\geqslant(2^{22}n)^{4t}\geqslant 2^{22}$, we have $3m^{-1/3}\leqslant 1/2e$, which implies
$\delta\leqslant 2\sqrt{\delta_{0}}+3m^{-1/3}\leqslant e^{-n/2t^{2}+1}+1/2e\leqslant 1/e^{2}+1/2e\leqslant 1/e.$
The second to last inequality is due to $n/2t^{2}+1\leqslant-2$. ∎
The bound $\delta\leqslant 1/e$ is clearly not sufficient for our purposes, but it can be easily upgraded using the Kruskal–Katona theorem, as the following claim shows.
Claim 41.
In terms and conditions of Claim 40 let $T_{0}\in\widetilde{\mathcal{T}}$ and $\delta<1/e$ be such that
$\big|\mathcal{F}\setminus(\mathcal{F}^{*}_{T_{0}}\vee\{T_{0}\})\big|=(1-\delta)m^{n-t}.$
Then the sum of sizes of $\mathcal{F}_{T}^{*},T\in\widetilde{T}\setminus\{T_{0}\}$, can be bounded as
$\sum_{T\in\widetilde{T}\setminus\{T_{0}\}}|\mathcal{F}_{T}^{*}|\leqslant\delta/2\cdot m^{n-t}$
Proof.
Let $(Z_{0},x_{0})$ be $T_{0}$ written in the restriction form. By averaging (see Claim 16), for each $T=(Z_{T},x_{T})\in\widetilde{\mathcal{T}}\setminus\{T_{0}\}$, find a restriction $y_{T}\in[m]^{Z_{0}\setminus Z_{T}}$ such that for $\mathcal{G}_{T}=\mathcal{F}^{*}_{T}(Z_{0}\setminus Z_{T}\to y_{T})$ we have $\mu_{[n]\setminus Z_{0}\cup Z_{T}}(\mathcal{G}_{T})\geqslant\mu^{-T}(\mathcal{F}^{*}_{T}).$ Theorem 23 implies
$|\partial_{t-1}\mathcal{F}_{T_{0}}^{*}|\geqslant(1-\delta)^{\frac{t-1}{n-t}}|\partial_{t-1}[m]^{[n]\setminus Z_{0}}|$
Families $\partial_{t-1}\mathcal{G}_{T}\subset\partial_{t-1}[m]^{[n]\setminus T_{0}\cup T}$ are disjoint for different $T\subset\widetilde{\mathcal{T}}\setminus\{T_{0}\}$ and also disjoint from $\partial_{t-1}\mathcal{F}^{*}_{T_{0}}\in\partial_{t-1}[m]^{[n]\setminus T_{0}}$ by Claim 35. Therefore the following inequality holds:
$\sum_{T\in\widetilde{\mathcal{T}}\setminus\{T_{0}\}}|\partial_{t-1}\mathcal{G}_{T}|\leqslant\left(1-(1-\delta)^{\frac{t-1}{n-t}}\right)\cdot|\partial_{t-1}[m]^{[n]\setminus Z_{0}}|\leqslant$
$\leqslant\frac{t-1}{n-t}\cdot\delta|\partial_{t-1}[m]^{[n]\setminus Z_{0}}|<\delta|\partial_{t-1}[m]^{[n]\setminus Z_{0}}|.$ (13)
We now use a contraposition of Theorem 23, i.e. Corollary 24, to bound the measure of a family via the size of its shadow. For any $T\in\widetilde{\mathcal{T}}\setminus\{T_{0}\}$, we have
$\mu^{-(Z\cup Z_{0})}(\mathcal{G}_{T})\leqslant\left(\frac{|\partial_{t-1}\mathcal{G}_{T}|}{|\partial_{t-1}[m]^{[n]\setminus(Z_{T}\cup Z_{0})}|}\right)^{(n-t-|Z_{0}\setminus Z_{T}|)/(t-1)}\leqslant$
$\leqslant\left(\frac{|\partial_{t-1}\mathcal{G}_{T}|}{|\partial_{t-1}[m]^{[n]\setminus Z_{0}}|\cdot{n-t-|Z_{0}\setminus Z_{T}|\choose t-1}/{n-t\choose t-1}}\right)^{(n-2t)/(t-1)}.$
Previously, we chose the restrictions $y_{T}$ in such a way that the measure of the remaining part was at least the measure of the original family. It allows us to bound the sum of measures of $\mathcal{F}_{T}^{*}$ as follows:
$\sum_{T\in\widetilde{\mathcal{T}}\setminus\{T_{0}\}}\mu_{[n]\setminus Z_{T}}(\mathcal{F}^{*}_{T})\leqslant\sum_{T\in\widetilde{\mathcal{T}}\setminus\{T_{0}\}}\mu_{[n]\setminus Z_{T}\cup Z_{0}}(\mathcal{G}_{T})$
$\qquad\leqslant\max_{T\in\widetilde{\mathcal{T}}\setminus\{T_{0}\}}\left(\frac{{n-t\choose t-1}}{{n-t-|Z_{0}\setminus Z_{T}|\choose t-1}}\right)^{\frac{n-2t}{t-1}}\cdot\sum_{T\in\widetilde{\mathcal{T}}\setminus\{T_{0}\}}\left(\frac{|\partial_{t-1}\mathcal{G}_{i}|}{|\partial_{t-1}[m]^{[n]\setminus Z_{0}}|}\right)^{\frac{n-2t}{t-1}}.$ (14)
We analyze these two factors separately. We have
$\max_{T\in\widetilde{\mathcal{T}}\setminus\{T_{0}\}}\left(\frac{{n-t\choose t-1}}{{n-t-|Z_{0}\setminus Z_{T}|\choose t-1}}\right)^{\frac{n-2t}{t-1}}$
$\qquad\leqslant\max_{T\in\widetilde{\mathcal{T}}\setminus\{T_{0}\}}\left(\frac{(n-t)(n-t+1)\cdot\ldots\cdot(n-2t+2)}{(n-t-|Z_{0}\setminus Z_{T}|)\cdot\ldots\cdot(n-2t-|Z_{0}\setminus Z_{T}|+2)}\right)^{\frac{n-2t}{t-1}}$
$\qquad\leqslant\max_{T\in\widetilde{\mathcal{T}}\setminus\{T_{0}\}}\left(1+\frac{|Z_{0}\setminus Z_{T}|}{n-2t-|Z_{0}\setminus Z_{T}|+2}\right)^{(t-1)\frac{n-2t}{t-1}}\leqslant\left(1+\frac{t}{n-3t}\right)^{n-2t}$
$\qquad\leqslant\left(1+\frac{2t}{n}\right)^{n}\leqslant e^{2t}.$ (15)
To analyze the second factor, note that the function $x\mapsto x^{\frac{n-2t}{t-1}}$ is convex, thus, $(x_{1},\ldots,x_{k})\to\sum_{i=1}^{k}{x_{i}^{\frac{n-2t}{t-1}}}$ is also convex. Put
$u_{T}=|\partial_{t-1}\mathcal{G}_{T}|/|\partial_{t-1}[m]^{[n]\setminus Z_{0}}|.$
Inequality (13) implies $\sum_{T\in\widetilde{\mathcal{T}}\setminus\{T_{0}\}}u_{T}<\delta$, thus, we can bound the second factor of (14) by a convex program
$\max_{u_{T},T\in\widetilde{\mathcal{T}}\setminus\{T_{0}\}}\sum_{T\in\widetilde{T}\setminus\{T_{0}\}}u_{T}^{\frac{n-2t}{t-1}}$
$\qquad\text{subject to}\quad\begin{cases}u_{T}\geqslant 0,\\ \sum_{T\in\widetilde{T}\setminus\{T_{0}\}}u_{T}\leqslant\delta.\end{cases}$
A convex function attains its maximum over a simplex at one of its vertices, which implies
$\sum_{T\in\widetilde{\mathcal{T}}\setminus\{T_{0}\}}\left(\frac{|\partial_{t-1}\mathcal{G}_{i}|}{|\partial_{t-1}[m]^{[n]\setminus Z_{0}}|}\right)^{\frac{n-2t}{t-1}}\leqslant\left(\frac{\delta|\partial_{t-1}[m]^{[n]\setminus Z_{0}}|}{|\partial_{t-1}[m]^{[n]\setminus Z_{0}}|}\right)^{\frac{n-2t}{t-1}}\leqslant\delta^{\frac{n-2t}{t-1}}.$ (16)
Combining (14), (15), (16) and the fact that $\delta<1/e$ we obtain
$\sum_{T\in\widetilde{\mathcal{T}}\setminus{T_{0}}}|\mathcal{F}_{T}^{*}|\leqslant m^{n-t}\cdot\sum_{T\in\widetilde{\mathcal{T}}\setminus\{T_{0}\}}\mu_{[n]\setminus Z_{T}}(\mathcal{F}^{*}_{T})\leqslant m^{n-t}\cdot e^{2t}\cdot\delta^{\frac{n-2t}{t-1}}\leqslant$
$\leqslant m^{n-t}\cdot\delta^{\frac{n-2t}{t-1}-2t}\leqslant m^{n-t}\delta^{2}\leqslant\delta/2\cdot m^{n-t}.$
The second to last inequality holds because $\frac{n-2t}{t-1}-2t\geqslant 2$. ∎
Now we can get a better bound on $\delta$ using the previous claim.
Claim 42.
In terms and conditions of Claim 40, there are $T_{0}\in\widetilde{\mathcal{T}}$ and $\delta<6m^{-1/3}$ such that
$\big|\mathcal{F}^{*}_{T_{0}}\vee\{T_{0}\}\big|=(1-\delta)m^{n-t}.$
Proof.
According to Claim 40 and Claim 41, there are $T_{0}\in\widetilde{\mathcal{T}}$ and $\delta\leqslant 1/e$ such that
$\big|\mathcal{F}^{*}_{T_{0}}\vee\{T_{0}\}\big|=(1-\delta)m^{n-t}$
and
$\sum_{T\in\widetilde{\mathcal{T}}\setminus\{T_{0}\}}|\mathcal{F}_{T}^{*}|\leqslant\delta/2\cdot m^{n-t}.$
The decomposition (12) together with $|\mathcal{F}|\geqslant m^{n-t}$ implies
[…]
Therefore $\delta<6m^{-1/3}$. ∎
6.4 Proof of Theorem 33
Proof of Theorem 33.
Let $\mathcal{F}\subset[m]^{n}$ be the maximal (t1)(t-1)-avoiding family, $|\mathcal{F}|\geqslant m^{n-t}$. We apply Lemma 14 to $\mathcal{F}$ with $\tau=m^{1/12t},q=24t^{2}$ and obtain a decomposition
$\mathcal{F}=\mathcal{R}\cup\bigcup_{i\in[l]}\mathcal{F}_{i},$
such that for all $i\in[l]$ there are some $Z_{i}\subset[n],|Z_{i}|\leqslant q$ and $x_{i}\in[m]^{Z_{i}}$ for which we have $\mathcal{F}_{i}\subset\mathcal{F}[Z_{i}\rightarrow x_{i}]$. Besides, $\mu(\mathcal{R})\leqslant\tau^{-q}$, and, for all $i$, we have $\mu_{Z_{i}\rightarrow x_{i}}(\mathcal{F}_{i}(Z_{i}\rightarrow x_{i}))\geqslant\tau^{-q}$ and $\mathcal{F}_{i}(Z_{i}\rightarrow x_{i})$ is $\tau$-homogenous with respect to $\mu_{Z_{i}\rightarrow x_{i}}$
Put $\mathcal{F}^{\prime}=\bigcup_{i\in[l]}\mathcal{F}_{i}$, $\mathcal{S}=\{(Z_{1},x_{1}),\ldots,(Z_{l},x_{l})\}$. Now we apply Lemma 21 to $\mathcal{F}^{\prime}$ playing the role of $\mathcal{F}$ and $\mathcal{S}$ being its $\tau$-homogenous core. We need to verify that all quantitative conditions hold. First, $t\leqslant n+1$ and $n\geqslant q+t-1=24t^{2}+t-1$ hold from the statement of Theorem 33. Second, $q=24t^{2}\geqslant t$ because $t\geqslant 1$. Finally, for $\alpha=(5n)^{-1}\in[0,(4n)^{-1})$ we have
$m\geqslant m^{1/12}\cdot(2^{22}n)^{t}\geqslant\max\{2^{15}\lceil\log_{2}n\rceil,48t^{2}\}\cdot 2\cdot(5n)^{t-1}\cdot m^{1/12}=$
$\max\{2^{15}\lceil\log_{2}n\rceil,2q\}\cdot 2\alpha^{1-t}\tau^{t}.$
The second inequality is due to $2^{22}n\geqslant 2\cdot 2^{15}\lceil\log_{2}n\rceil$ and $2^{22}n\geqslant 2^{22}\cdot 50t^{2}\geqslant 2\cdot 48t^{2}$.
Thus, we obtain families $\mathcal{U}_{1},\ldots,\mathcal{U}_{l}$ satisfying (1 — 3) of Lemma 21. It is easy to see that $\mathcal{U}=\bigcup_{i\in[l]}\mathcal{U}_{i}$ is an $(\mathcal{S},2,t)$-system. We now apply Corollary 34 to $\mathcal{U}$ and obtain a set of restrictions $\widetilde{\mathcal{T}}$, satisfying (9) and (10). We can decompose $\mathcal{F}$ as
$\mathcal{F}=\mathcal{R}\cup\mathcal{F}[\widetilde{\mathcal{T}}]\cup\bigcup_{S\in\mathcal{S}\setminus\mathcal{S}[\widetilde{\mathcal{T}}]}\mathcal{F}_{S}.$
Then we have
$|\mathcal{F}\setminus\mathcal{F}[\widetilde{\mathcal{T}}]|\leqslant|\mathcal{R}|+\sum_{S\in\mathcal{S}\setminus\mathcal{S}[\widetilde{\mathcal{T}}]}|\mathcal{F}_{S}|\leqslant m^{n}\cdot\tau^{-q}+(1-2\alpha n)^{-1}\sum_{S\in\mathcal{S}\setminus\mathcal{S}[\widetilde{\mathcal{T}}]}|\mathcal{U}_{S}|\leqslant$
[…]
We partition $\mathcal{F}[\widetilde{\mathcal{T}}$] into families $\mathcal{F}_{T}\subset\mathcal{F}[Z_{T}\to x_{T}]$ for all $(Z_{T},x_{T})=T\in\widetilde{\mathcal{T}}$, i. e.
$\mathcal{F}[\widetilde{\mathcal{T}}]=\bigsqcup_{T\in\widetilde{\mathcal{T}}}\mathcal{F}_{T}.$
Note that $n,m,t,\mathcal{F},\mathcal{R},\widetilde{\mathcal{T}}$ meet conditions of Claim 42. Thus, due to this claim, there are $(Z_{0},x_{0})=T_{0}\in\widetilde{\mathcal{T}}$ and $\delta<6m^{-1/3}$ such that
$\big|\mathcal{F}^{*}_{T_{0}}\vee\{T_{0}\}\big|=(1-\delta)m^{n-t}.$
It remains to show that in fact $\delta=0$ and $\mathcal{F}=[m]^{n}[Z_{0}\to x_{0}]=[m]^{n}[T_{0}]$. Put $\mathcal{R}=\mathcal{F}\setminus\mathcal{F}[T_{0}]$. To bound $|\mathcal{R}|$ we need the following claim.
Claim 43.
The families $\partial_{t-1}\mathcal{F}_{T_{0}}^{*}$ and $\partial_{t-1}\left(\mathcal{R}|_{[n]\setminus Z_{0}}\right)$ do not intersect.
Proof.
Assume they do, $T=(Z,x)$ is a common element, |Z|=t1|Z|=t-1. Fix arbitrary $F\in\mathcal{F}_{T_{0}}^{*}[Z\to x],G\in\mathcal{R}[Z\to x].$ Put $p=\operatorname{agr}(x_{0},G|_{Z_{0}})$. Since no set from $\mathcal{R}$ contains $T_{0}$, we have p<t1p<t-1.
Fix $W\subset Z,|W|=t-1-p.$ Let $Y$ be $F|_{W}$, note that $|Y|\leqslant t-1$. By the definition of $\mathcal{F}_{T_{0}}^{*}$ the covering number of $\mathcal{F}^{*}_{T_{0}}(T_{0}\cup Y)$ is at least $n+1$ thus there is some $F_{1}\in\mathcal{F}_{T_{0}}(T_{0})[Y]$ that avoids $G$ outside of $Y$. Then, we have $\operatorname{agr}(F_{1}\sqcup T_{0},G)=t-1$, a contradiction. ∎
Using Theorem 23, we get that
$|\partial_{t-1}\mathcal{F}_{T_{0}}^{*}|\geqslant(1-\delta)^{\frac{t-1}{n-t}}|\partial_{t-1}[m]^{n\setminus Z_{0}}|.$
According to the previous claim, using disjointness,
$|\partial_{t-1}\left(\mathcal{R}_{[n]\setminus Z_{0}}\right)|\leqslant\left(1-(1-\delta)^{\frac{t-1}{n-t}}\right)|\partial_{t-1}[m]^{n\setminus Z_{0}}|\leqslant\frac{t-1}{n-t}\delta|\partial_{t-1}[m]^{n\setminus Z_{0}}|.$
Finally, using Corollary 24 once again we get
$|\mathcal{R}|\leqslant m^{t}\cdot\left|\partial_{n-t}\left(\mathcal{R}_{[n]\setminus Z_{0}}\right)\right|\leqslant m^{t}\cdot\left(\frac{\delta(t-1)}{n-t}\right)^{\frac{n-t}{t-1}}m^{n-t}\leqslant$
$\leqslant m^{n}\cdot m^{-1/3\cdot\frac{n-t}{t-1}}<m^{n-3t}.$
The last inequality is true because $\frac{1}{3}\cdot\frac{n-t}{t-1}<3t$. Finally we apply Lemma 31 to $\mathcal{F}$, $Z_{0}$ playing the role of $Z$ and $x_{0}$ playing the role of $x$, which implies $\mathcal{F}=[m]^{n}[T_{0}]$.
7 Proof of the main theorem
Proof of Theorem 3.
Let $\widehat{n}_{0}(t),\widehat{m}_{0}(t)$ be polynomials from Corollary 32 and set
$m_{0}(t)=\widehat{m}_{0}(t),\quad\bar{n}_{0}(t)=\left(88\ln 2+4\right)t\cdot\widehat{n}_{0}(t)+50t^{2}.$
Consider integers $n,m,t$ and a family $\mathcal{F}\subset[m]^{n}$ such that $\operatorname{agr}(x,y)\neq t-1$ for any $x,y\in\mathcal{F}$ and $|\mathcal{F}|\geqslant m^{n-t}$. We will prove that if $m\geqslant m_{0}(t)$ and $n\geqslant\bar{n}_{0}(t)\ln n$, then we may apply either Corollary 32 or Theorem 33 and conclude $\mathcal{F}=[m]^{n}[Z\rightarrow x]$ for some $Z\subset n$ such that $|Z|=t$ and $x\in[m]^{Z}$.
If $n\geqslant\widehat{n}_{0}(t)\ln m$, we apply Corollary 32 and get the desired conclusion. Otherwise, we get
$m\geqslant e^{\frac{n}{\widehat{n}_{0}(t)}}\geqslant e^{(88\ln 2+4)t\ln n}\geqslant e^{(88\ln 2)t+4t\ln n}=\left(2^{22}n\right)^{4t}.$
Also we have $n\geqslant\bar{n}_{0}(t)\geqslant 50t^{2}$. Thus, we may apply Theorem 33 and get $\mathcal{F}=[m]^{n}[Z\rightarrow x]$ for some $Z\subset n$ such that $|Z|=t$ and $x\in[m]^{Z}$.
Finally, we will show that for $n\geqslant\widetilde{n}_{0}(t)=2^{32}t^{3}\ln t$, we have $n\geqslant\bar{n}_{0}(t)\ln n$. Indeed, if $n\geqslant\widetilde{n}_{0}(t)$, we have
$n$ $=n^{1-\frac{1}{3\ln t}}n^{1/3\ln t}\geqslant{\widetilde{n}_{0}(t)}^{1-1/3\ln t}\cdot\frac{\ln n}{3\ln t}\geqslant 2^{16}t^{3}\ln t\cdot(t^{3}\ln t)^{-\frac{1}{3\ln t}}\cdot\frac{\ln n}{3\ln t}$
$\geqslant 2^{16}e^{-4/3}/3\cdot t^{3}\ln n\geqslant\bar{n}_{0}(t)\ln n.$
In particular, choosing $n_{0}(t)=2^{32}t^{4}$, we get the result. ∎
Acknowledgements
We thank Andrey Kupavskii for his participation in the initial discussions and for his fruitful suggestions on the final draft of this paper. We are also grateful to Noam Lifshitz for his comments on Theorem 10. Additionally, Fedor Noskov thanks Pierre Youssef for introducing the semigroup method and its widespread applications.
Appendix A Proof of Lemma 12
Lemma 12 is a quantitative analogue of Theorem 2.3 from [42] and may be obtained by a verbatim repetition of its proof. We will prove the lemma by contradiction in the following way:
Step 1. Suppose that $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$ are cross-agreeing. Using a certain transformation, we will transform the families $\mathcal{F}_{i}$ into $\widetilde{\mathcal{F}}_{i}$ in such a way that they remain cross-agreeing, become monotone, and have large $p$-biased measure.
Step 2. Next, we will move from the $p$-biased measure to the uniform measure $\mu_{1/2}$ on $\{0,1\}^{n}$, and use the monotonicity of $\widetilde{\mathcal{F}}_{i}$ to show that $\mu_{1/2}(\widetilde{\mathcal{F}}_{i})$ is sufficiently large.
Step 3. Finally, we will show that $\mu_{1/2}(\widetilde{\mathcal{F}}_{1})+\mu_{1/2}(\widetilde{\mathcal{F}}_{2})>1$, which leads to a contradiction with cross-agreeing.
One part of the transformation in Step 1 will be the following compression operator, defined in Section 4.2 of [42].
Definition 8.
For any $i\in[n]$ and $j\in[m]$, we define the compression operator $T_{i,j}:[m]^{n}\to[m]^{n}$. For $x\in[m]^{n}$, set $T_{i,j}(x)=y\in[m]^{n}$ where $y_{r}=x_{r}$ for all $r\neq i$, and $y_{i}=x_{i}$ if $x_{i}\neq j$, or $y_{i}=1$ if $x_{i}=j$.
We also define a compression operator, also denoted $T_{i,j}$, on codes:
$T_{i,j}(\mathcal{F})=\{x\;:\;T_{i,j}(x)\in\mathcal{F}\}\cup\{T_{i,j}(x)\;:\;x\in\mathcal{F}\}.$
We define $T_{i}=T_{i,2}\circ T_{i,3}\circ\cdots\circ T_{i,m}$ for any $i\in[n]$, and $T=T_{1}\circ T_{2}\circ\cdots\circ T_{n}$.
We call $\mathcal{F}\subset[m]^{n}$ compressed if $T(\mathcal{F})=\mathcal{F}$.
Another part of the transformation in Step 1 will be the transformation from compressed codes in $m]^{n}$ to monotone families in the cube $\{0,1\}^{n}$.
Definition 9 (Definition 4.2 from [42]).
We define $h:[m]\to\{0,1\}$ by $h(1)=1$ and $h(a)=0$ for all $a\neq 1$, and $h^{\otimes n}:[m]^{n}\to\{0,1\}^{n}$ by $h^{\otimes n}(x)=\big(h(x_{1}),\ldots,h(x_{n})\big)$.
To describe the properties of these transformations we need the following definition for $\{0,1\}^{n}$, which is stricter than that for $m]^{n}$.
Definition 10.
Two families $\mathcal{F},\mathcal{G}\subset\{0,1\}^{n}$ are called cross-agreeing if
$\forall x\in\mathcal{F},\forall y\in\mathcal{G},\quad\left|\{i\in[n]:x_{i}=y_{i}=1\}\right|\geqslant 1.$
Definition 11.
A family $\mathcal{A}\subset\{0,1\}^{n}$ is called monotone if, for any $x,y\in\{0,1\}^{n}$ such that $x\leqslant y$ coordinatewise, $x\in\mathcal{A}$ implies $y\in\mathcal{A}$.
The transformation at Step 1 will have the following properties, that are direct consequences of Fact 4.1 and Fact 4.3 from [42]
Lemma 44.
Let $\mathcal{F}_{1},\mathcal{F}_{2}\subset[m]^{n}$ are cross-agreeing. Let $\widetilde{\mathcal{F}}_{i}=h^{\otimes n}(T(\mathcal{F}_{i}))$. Then
  1. 1.
    For each $i$ the family $\widetilde{\mathcal{F}}_{i}$ is monotone and $\mu_{1/m}(\widetilde{\mathcal{F}}_{i})\geqslant\mu(\mathcal{F}_{i})$.
  2. 2.
    $\widetilde{\mathcal{F}}_{1}$ and $\widetilde{\mathcal{F}}_{2}$ are cross-agreeing.
For Step 2, we will need the following lemma of Ellis, Keller, and Lifshitz [13] to obtain a lower bound on $\mu_{1/2}$ from $\mu_{1/m}$.
Lemma 45.
Suppose $0\leqslant p\leqslant q\leqslant 1$, $\alpha\geqslant 0$ and $\mathcal{F}\subset\{0,1\}^{n}$ is monotone. If $\mu_{p}(\mathcal{F})\geqslant p^{\alpha}$, then $\mu_{q}(\mathcal{F})\geqslant q^{\alpha}$.
For Step 3, we will need the following result.
Lemma 46 (Fact 4.11 from [42]).
If $\mathcal{A},\mathcal{B}\subset\{0,1\}^{n}$ are cross-agreeing, then
$\mu(\mathcal{A})+\mu(\mathcal{B})\leqslant 1.$
Now we are ready to prove Lemma 12.
Proof.
Step 1. Assume, for the sake of contradiction, that the families $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$ are cross-agreeing. Define the transformed families:
$\widetilde{\mathcal{F}}_{i}=h^{\otimes n}(T(\mathcal{F}_{i})),\quad\text{for }i=1,2.$
By Lemma 44 the families $\widetilde{\mathcal{F}}_{1}$ and $\widetilde{\mathcal{F}}_{2}$ are cross-agreeing, monotone and satisfy
$\mu_{1/m}(\widetilde{\mathcal{F}}_{i})\geqslant\mu(T(\mathcal{F}_{i}))=\mu(\mathcal{F}_{i}),\quad\text{for }i=1,2,$
where $\mu_{p}$ denotes the $p$-biased measure on $\{0,1\}^{n}$.
Step 2. Now consider $\widetilde{\mathcal{F}}_{i}$ under the uniform measure $\mu_{1/2}$. By monotonicity, we have $\mu_{1/2}(\widetilde{\mathcal{F}}_{1})\geqslant\mu_{1/m}(\widetilde{\mathcal{F}}_{1})$. By Lemma 45, we obtain
$\mu_{1/2}(\widetilde{\mathcal{F}}_{2})\geqslant\left(\frac{1}{2}\right)^{\log_{1/m}\left(\mu_{1/m}(\widetilde{\mathcal{F}}_{2})\right)}=\mu_{1/m}(\widetilde{\mathcal{F}}_{2})^{\log_{m}(2)}.$
Step 3. Combining the above inequalities yields
$\mu_{1/2}(\widetilde{\mathcal{F}}_{1})+\mu_{1/2}(\widetilde{\mathcal{F}}_{2})$ $\geqslant\mu_{1/m}(\widetilde{\mathcal{F}}_{1})+\mu_{1/m}(\widetilde{\mathcal{F}}_{2})^{\log_{m}(2)}$
$\geqslant\mu(\mathcal{F}_{1})+\mu(\mathcal{F}_{2})^{\log_{m}(2)}$
$>1,$
which contradicts Lemma 46, completing the proof. ∎